### v1

# A Large-scale Empirical Study on Large Language Models for Election Prediction

Abstract
Can Large Language Models (LLMs) accurately predict election outcomes? While LLMs have demonstrated impressive performance in healthcare, legal analysis, and creative applications, their capabilities in election forecasting remain uncertain. Notably, election prediction poses unique challenges: limited voter-level data, evolving political contexts, and the complexity of modeling human behavior. In the first part of this paper, we explore and introduce a multi-step reasoning framework for election prediction, which systematically integrates demographic, ideological, and time-sensitive factors. Validated on 2016 and 2020 real-world data and extensive synthetic personas, our approach adapts to changing political landscapes, reducing bias and significantly improving predictive accuracy. We further apply our pipeline to the 2024 U.S. presidential election, illustrating its ability to generalize beyond observed historical data. Beyond enhancing accuracy, the second part of the paper provides insights into the broader implications of LLM-based election forecasting. We identify potential political biases embedded in pretrained corpora, examine how demographic patterns can become exaggerated, and suggest strategies for mitigating these issues. Together, this project, a large-scale LLM empirical study, advances the accuracy of election predictions and establishes directions for more balanced, transparent, and context-aware modeling in political science research and practice.
A Large-scale Empirical Study on Large Language Models for Election Prediction
Chenxiao Yu1, Zhaotian Weng1, Yuangang Li1, Zheng Li2, Xiyang Hu3, Yue Zhao1 1University of Southern California  2Arima  3Carnegie Mellon University {cyu96374, wengzhao, yuangang, yzhao010}@usc.edu, winston@arimadata.com, xiyanghu@cmu.edu Corresponding authors
1 Introduction
Large Language Models (LLMs) have shown strong capabilities across diverse domains, including drug discovery Liu et al. (2024), legal analysis Chalkidis et al. (2022), and creative tasks Yang et al. (2022). Their strength lies in not only understanding and generating text but also leveraging vast common knowledge Roberts et al. (2020), simulating diverse personas Hu and Collier (2024), and modeling human behavior in complex social contexts Bommasani et al. (2021). LLMs excel in human-like reasoning Zhou et al. (2020); AlKhamissi et al. (2022) and have been applied to simulate decision-making in various settings Zhou et al. (2023); Ziems et al. (2024).
Notably, recent work has extended LLMs to political science, analyzing policy documents, campaign speeches, and public sentiment Xu (2022); Haq et al. (2023); Li et al. (2024b). Beyond language processing, their ability to integrate knowledge and apply human-like reasoning positions them as powerful tools for simulating the intricate dynamics of political decision-making Argyle et al. (2023); Bisbee et al. (2024).
Figure 1: Demonstration of three prompt designs in §2.2. V1 is the direct prompt on voter demographic information, while V2 introduces time-dependent information to capture candidates’ agenda and V3 also uses multi-step reasoning. In this example for 2020 Ohio result prediction, only V3 can accurately predict the results, demonstrating the importance of leveraging both time-dependent information and multi-step reasoning for election result prediction.
Background. Despite LLMs’ achievements in more straightforward political science tasks, their capacity to handle more challenging efforts, such as election prediction, is not well understood Lerer et al. (2022). LLMs’ access to extensive historical data suggests potential for forecasting, but election prediction introduces unique challenges. First, the high cost of acquiring voter-level data complicates both experimentation and model validation. Second, election forecasting involves modeling not only voter behavior, but also the shifting political context, making it unclear whether text-based data alone can capture the full complexity Graefe (2014). Third, accurate predictions require reasoning about multiple factorsHolbrook (2016). Whether LLMs can perform this more advanced reasoning remains unknown. Wei et al. (2022).
This Work. We present a large-scale empirical study to address two key questions in election prediction with LLMs: (i) how can LLMs be leveraged to achieve accurate election predictions, and (ii) beyond accuracy, what additional considerations are crucial for understanding LLM capabilities in election forecasting tasks?
Contribution 1: Multi-Step Reasoning Framework for Accurate Election Prediction (§2). To handle the scarcity of detailed voter-level data, we employ the Sync synthetic data generation framework Li et al. (2020b), which probabilistically reconstructs individual demographic and behavioral profiles from aggregated public datasets. We pair this synthetic data with real-world datasets, notably the American National Election Studies (ANES) Time Series Studies (2022), ensuring that our model aligns with actual voting behaviors. Our approach also adapts to evolving political conditions by incorporating time-dependent factors. We aggregate information from presidential campaign data, including candidates’ policy agendas and backgrounds, to align the model with changing political environments Holbrook (2016). Finally, we introduce a multi-step reasoning framework guided by Chain of Thought prompting Wei et al. (2022). By breaking down the prediction process into intermediate steps, the model can systematically integrate demographic details, ideological alignment, and time-sensitive information. This approach strengthens the model’s accuracy and mitigates biases and overfitting issues observed in simpler models. As shown in Figs. 1 and 2, we refine our pipeline through iterations. The final version significantly improves predictive accuracy and alignment with real-world data via extensive experiments.
Contribution 2: New Insights and Directions Beyond Accurate Prediction (§3). Our experiments uncover several key insights that extend beyond raw predictive accuracy. First, while our multi-step reasoning pipeline reduces bias and improves alignment with real-world outcomes, residual skewness persists, highlighting the influence of pretrained corpora on political predictions. Second, although LLMs can replicate known demographic voting patterns, they tend to exaggerate these differences, raising concerns about reinforcing stereotypes. Finally, our results indicate that integrating richer contextual information and refining prompt strategies can further enhance predictive reliability. We thus propose meaningful future directions on reducing political skewness in pretrained corpora and mitigating stereotype biases in election prediction.
2 Part 1: Accurate Election Prediction via a Multi-Step LLM Pipeline
How can we effectively leverage LLMs to predict election results? In this work, we simulate each voter’s decision-making process by providing LLMs with detailed voter information and asking them to predict voter preferences. To achieve this, we focus on two key aspects: (1) establishing an evaluation framework with appropriate datasets that contain voter-level information, and (2) designing an LLM-based pipeline for accurate election predictions. In §2.1, we describe the datasets and the evaluation methods. We then provide an overview of our design approach in §2.2 with three progressive pipelines. Finally, we evaluate the performance of the pipelines on datasets in §2.3.
Figure 2: Progressive design of LLM pipelines for election predictions. V1: Demographic-only Prompting2.2.1) uses static demographic personas but lacks temporal context. V2: Time-based Prompting2.2.2) adds election-year policy shifts and candidate information, but the single-step prompt setup leads to overloaded inputs that reduce prediction accuracy. V3: Multi-step Reasoning2.2.3) divides the decision-making process into multiple steps, improving reasoning quality and producing results that closely match real-world outcomes. All versions aggregate individual predictions into state-level results to capture broader election trends.
2.1 Datasets, Evaluation, and Settings
Datasets. This study leverages two primary data sources: (1) Public Benchmarks: The American National Election Studies (ANES) 2016 and 2020 Time Series data Studies (2019, 2022), which provide comprehensive demographic and political ideology information for real respondents. This dataset is a benchmark to evaluate how accurately LLMs simulate voter-level behavior in alignment with real-world data. (2) For state-level voting simulations, we use voter-level synthetic data generated via advanced ML techniques based on aggregated information Li et al. (2020b). Personas are randomly sampled for each state at specified ratios, and predictions are compared with actual election outcomes from 2020 Federal Election Commission (2021) and 2024 NBC News (2024).
Both datasets contain non-personally identifiable [^1] This project has been reviewed by the IRB and considered exempt as the datasets are non-personally identifiable., voter-level information. Detailed partition and sampling methodologies are in Appx. B.1.
Evaluation Method. To evaluate performance on public benchmarks and state-level simulations, we firstly assess how closely LLM simulations align with actual voting results. The calculation follows the approach outlined in Xie et al. (2024); Argyle et al. (2023) and is defined as:
Predicted Proportion Ratio $P(s)$
$P(s)=\frac{\text{Republican Votes}}{\text{Republican Votes}+\text{Democratic Votes}}$ (1)
Here, $s$ represents the unit of analysis: it can refer to an individual persona (as in public benchmarks like ANES) or an entire state (as in state-level simulations). This ratio $P(s)$ measures the proportion of votes predicted for the Republican Party relative to the total votes for the two major parties, excluding those with no preference.
Hardware and LLM Settings. Our experiments were conducted on a GPU server with an AMD EPYC Milan 7763 processor with 6 NVIDIA RTX A6000 Ada GPUs. For the LLM, we use OpenAI’s GPT-4o model for election predictions. Meta’s LLaMA 3.1 (405B) model is employed in intermediate steps to provide neutral summarizations of time-dependent information, enhancing certain pipelines for capturing temporal dynamicsFeng et al. (2023).
2.2 Our Progressive Design of LLM Pipelines
In this section, we present our progressive design for generating voter-level election predictions using LLMs. As shown in Fig. 2, we develop three versions of the pipeline. Each version addresses a key shortcoming of its predecessor and integrates more detailed information and reasoning processes.
  1. V1:
    Demographic-only Prompting (§2.2.1): This initial version uses static demographic personas to guide voter-level predictions. It is simple but does not capture changes in candidates’ focus over time.
  2. V2:
    Single-step Prompting with Time-based Information (§2.2.2): Here, we add election-year-specific details, such as policy agendas and candidate backgrounds, to make predictions more dynamic. However, packing all information into a single prompt can overwhelm the model, limiting its reasoning depth.
  3. V3:
    Multi-step Reasoning with Contextual Information (§2.2.3): This version separates the prediction process into a sequence of steps. By breaking down the reasoning, the model can integrate demographics, candidate profiles, and political context better, for predictions that better match real-world results.
2.2.1 Version 1: Demographic-only Prompting
This initial version directly prompts the LLM with a persona that includes demographic attributes (e.g., age, gender, income) and asks how that persona would vote Xie et al. (2024); Argyle et al. (2023).
Task: You are persona [age, gender, ethnicity, marital status, household size, presence of children, education level, occupation, individual income, family income, and place of residence.] The current year is [year].
Please answer the following question as if you were the resident: 1. As of today, will you vote for the Democratic Party (Joe Biden), the Republican Party (Donald Trump), or do you have no preference?
Options: Democratic, Republican, No Preference
Specifying the year as 2020 avoids confusion because the LLM’s knowledge ends in 2023. The listed voting options follow Pew Research Center’s 2014 Political Polarization and Typology Survey Pew Research Center (2014).
Limitations: This version cannot adapt predictions to different time periods. Without incorporating temporal changes in candidates’ agendas or public opinion, predictions remain static, limiting their usefulness for evolving electoral contexts.
2.2.2 Version 2: Single-step Prompting with Time-based Information
Accurately modeling elections requires accounting for macro-level factors and time-specific variations Gao et al. (2022b). To improve realism, we extend our pipeline by incorporating election-year data from Ballotpedia, a widely used platform that provides campaign agendas, key policy positions, and candidates’ biographical details.
Ensuring that this time-based information is conveyed neutrally is important, given documented political biases in LLMs Feng et al. (2023). We compared GPT-4o and LLaMA3-405B for summarizing these details, and found that LLaMA3-405B provided more balanced outputs. We then integrated these balanced summaries into the prompts:
Task: You are persona [demographics]. The current year is [year]. [Two parties’ policy agenda]. [Presidential candidates’ biographical and professional backgrounds].
Please answer the following question as if you were the resident: 1. As of today, will you vote for the Democratic Party (Joe Biden), the Republican Party (Donald Trump), or do you have no preference?
Options: Democratic, Republican, No Preference
Limitations:. Although adding time-dependent context makes predictions more dynamic, it introduced a skew towards Democratic candidates in both “deep red” states and key swing states. For example, predictions across five historically Republican states (e.g., Alabama, South Carolina) and multiple swing states (e.g., Texas, Florida) leaned more Democratic than expected. This pattern aligns with earlier observations that GPT-4o leans liberal Feng et al. (2023). Interestingly, when tested on the ANES 2020 dataset Studies (2022), which includes ideological self-placement, the skew was less pronounced. Additional analysis showed that removing ideological self-placement shifted outcomes significantly, suggesting that such ideological cues help correct biases introduced by purely demographic or time-based prompts.
2.2.3 Version 3: Multi-step Reasoning with Contextual Information
To address the limitations in Version 2, we introduce a multi-step prompting pipeline. Drawing on the Chain of Thought prompting strategy Wei et al. (2022), this approach decomposes the prediction process into intermediate steps, enabling more systematic reasoning and improved accuracy.
The procedure involves two main steps: (1) Conservative-Liberal Spectrum Placement: The model is given a persona and current policy positions of both parties, and is asked to position the persona on a conservative-liberal spectrum. (2) Extended Persona and Voting Simulation: This ideological self-placement is then incorporated into the persona. Together with time-based information, this enriched context is used in the second step to simulate voting behavior.
Step 1: You are a persona with [demographics]. The current year is [year]. [Two parties’ policy agenda]. When it comes to politics, would you describe yourself as: No answer Very liberal Somewhat liberal Closer to liberal Moderate Closer to conservative Somewhat conservative Very conservative Step 2: You are a persona with [demographics]. Your [conservative-liberal spectrum]. The current year is [year]. [Two parties’ policy agenda]. [Presidential candidates’ biographical and professional backgrounds].
Please answer the following question as if you were the resident: 1. As of today, will you vote for the Democratic Party (Joe Biden), the Republican Party (Donald Trump), or do you have no preference?
Options: Democratic, Republican, No Preference
Applying the multi-step method at the state level for the 2020 election, we evaluated its performance across a range of scenarios, including “deep red” states (e.g., Alabama), “deep blue” states (e.g., California), and 11 swing and tipping-point states (details in Appendix C). This method demonstrated significantly improved accuracy, with predicted vote distributions closely aligning with actual results across almost all Chosen states.
Given these accuracy gains and enhanced alignment with real-world data, we adopt Version 3 as our final pipeline for election prediction. By structuring the reasoning process into multiple steps, this approach effectively mitigates skewness and captures the complex dynamics of voter behavior across diverse state contexts.
2.3 Empirical Validation and Comparative Analysis of the Proposed Pipelines
2.3.1 Public ANES Benchmark Evaluation
We first evaluated the pipelines using the ANES 2016 and 2020 Time Series datasets Studies (2019, 2022) to assess their overall performance and to verify V3’s ability to generate the Conservative-Liberal Spectrum feature. For V1 (Demographic-only Prompting) and V2 (Single-step Prompting with Time-based Information), we excluded the ideological spectrum, relying solely on demographic data. In contrast, V3 employs multi-step reasoning, inspired by Chain of Thought prompting Wei et al. (2022), to incorporate ideological alignment alongside demographics and time-sensitive factors. While the 2020 dataset lacked certain demographic variables—requiring restoration of the spectrum—the 2016 dataset allowed a direct comparison between LLM-generated (3rd Pipeline_G) and restored (3rd Pipeline_featureBack) spectra.
Figure 3: Comparison of the three pipelines on ANES 2016 and 2020. The y-axis shows the predicted ratio (Eq. 1), with 0.5 indicating balance. The red baseline represents ground truth ratios from the ANES dataset.
As shown in Fig. 3, excluding the spectrum-skewed predictions toward the election winner, V1 and V2 favored Republicans in 2016 (63.25%, 66.38% support for Trump) and Democrats in 2020 (30.14%, 10.02%). By contrast, V3 closely matched the actual results (e.g., 48.38% vs. an actual 47.7% for Trump in 2016). Notably, the LLM-generated spectrum in V3 slightly outperformed the restored version, showing the model’s ability to produce meaningful ideological features. These highlight the importance of incorporating ideological alignment to improve accuracy and reduce bias.
2.3.2 2020 State-Level Evaluation
State-level simulations using synthetic data were conducted to compare the model predictions against the official 2020 election outcomes reported by the Federal Election Commission (FEC). We selected five red states, five blue states, and 11 swing or tipping-point states. The V1 pipeline (Demographic-only Prompt) displayed a consistent bias toward the Democratic Party, even in historically Republican states such as South Carolina (SC) and Ohio (OH), underscoring the limitations of relying solely on demographic features. Introducing time-dependent, election-year-specific information in V2 (Time-dependent Prompt) reduced this bias, but the predictions still skewed Democratic.
In contrast, the V3 pipeline (Multi-step Reasoning) demonstrated the highest accuracy and effectively mitigated bias, closely aligning its predictions with real outcomes. It correctly forecasted the results in 9 out of 11 swing states, with only minor misestimations in North Carolina (NC) and Arizona (AZ), and accurately captured voter behavior in deeply red and blue states. These findings highlight V3’s capability to model complex voter dynamics and closely track actual election trends. Detailed results are provided in Appx. C.
2.3.3 2024 State-Level Evaluation
Before the official 2024 results were available, we employed V3 to forecast outcomes across all 50 states. For consistency and cost-efficiency, V1 and V2 were evaluated on the same subset of states. This analysis tested V3’s predictive performance in an unseen future context and allowed for comparisons with previous research Zhang et al. (2024a).
V3 performed robustly in 2024, which accurately predicted outcomes in all traditional red and blue states (with the exception of Alaska) and successfully forecasted Trump’s wins in Arizona (AZ), Wisconsin (WI), Florida (FL), and Texas (TX), as well as Harris’s victories in New Hampshire (NH) and Minnesota (MN). Although V3 slightly mispredicted several swing states, these errors were significantly smaller than the pronounced Democratic bias exhibited by V1 and V2.
Overall, these evaluations confirm V3’s strong ability to simulate voter behavior, reduce political bias, and improve alignment with real-world election patterns. Further details on the 2024 predictions are presented in Appx. D.
3 Beyond Accuracy: Insights and Future Directions Beneath Results
In §2, we proposed a voting prediction pipeline (§2.2.3) that demonstrated high accuracy using the ratio metric Eq. (1). This section delves deeper by introducing additional evaluation metrics and analyzing the broader implications of the results.
Using these metrics, we revisit the 2020 and 2024 state-level predictions to evaluate pipelines’ overall accuracy and examine systematical biases toward political parties (§3.1). Finally, we analyze how LLM simulations align with real-world demographic trends and discuss strategies for mitigating bias and enhancing the robustness of future LLM-based election forecasting methods (§3.2).
Additional Evaluation Metrics. We consider additional three evaluation metrics to assess our predictions against actual election results: Weighted Absolute Error (WAE) (Eq. (A4)), Weighted Mean Squared Error (WMSE) (Eq. (A5)), and Bias Metric (BM) (Eq. (A6)). These metrics provide a comprehensive evaluation of model performance. WAE and WMSE measure the degree of alignment between predicted proportions and actual outcomes, with WMSE placing greater emphasis on larger errors due to its squared formulation. BM quantifies systematic biases, indicating whether the predictions consistently favor one party over the other. Together, these metrics offer a nuanced understanding of both the magnitude of prediction errors and the potential biases within the pipelines. See Appx. B.2 for details.
3.1 Result Analysis on Systematical Accuracy and Bias
To align our evaluation of the 2020 and 2024 pipelines with other studies, we selected 11 swing and tipping-point states as the scope for analysis.
Metric Vanilla (%) V2 (%) V3 (%)
Weighted Absolute Error 22.78 14.97 5.24
Weighted Mean Squared Error 5.46 2.34 0.37
Bias Metric (BM) -22.78 -14.97 0.34
Table 1: Evaluation metrics for the 2020 election simulation across three pipelines (Vanilla, V2, and V3). V3 achieves the lowest errors and minimal bias, showing substantial improvement over the simpler approaches.
Metric Vanilla (%) V2 (%) V3 (%) C (%)
Weighted Absolute Error 21.35 25.96 3.49 4.70
Weighted Mean Squared Error 4.83 6.89 0.22 0.30
Bias Metric (BM) -21.35 -25.96 -2.95 1.26
Table 2: Evaluation metrics for the 2024 election simulation comparing three pipeline versions (Vanilla, V2, V3) and one external comparison model (C, from Zhang et al. (2024a)). As with 2020, V3 achieves the most accurate results (lowest WAE and WMSE). Although BM remains slightly negative for V3, it is greatly reduced compared to the Vanilla and V2 pipelines, indicating a less biased prediction. The comparison model (C) performs competitively, but does not surpass V3.
Results Analysis of 2020. We begin our extended analysis by examining the model performances on the 2020 election. Table 1 presents the results for three pipelines: Vanilla (V1), V2, and V3. Here, WAE and WMSE serve as measures of how closely predicted vote proportions match the actual results, while the BM indicates the presence and direction of any systematic lean. For the 2020 simulation, the Vanilla pipeline shows substantial bias towards Democratic Party, as evidenced by its large negative BM value and correspondingly high error rates. Incorporating time-dependent information in V2 reduces both WAE and WMSE compared to Vanilla, but still a notable Democratic skew persists, indicated by a still-negative BM. In contrast, V3 significantly lowers both absolute and squared errors and nearly eliminates bias, achieving a BM close to zero. This improvement confirms that the multi-step reasoning approach employed by V3—incorporating ideological placement and time-sensitive data—effectively counters the model’s initial skew and enhances overall accuracy.
Results Analysis of 2024. We next turn to the 2024 predictions. Table 2 compares the same three pipelines with an additional external comparison model (C) from Zhang et al. (2024a). The pattern observed in 2020 holds true in 2024: Vanilla and V2 again exhibit marked bias toward the Democratic Party, reflected in large negative BM values. Although V3 still shows a slight Democratic lean (BM = -2.95), it is significantly less than in the simpler pipelines. V3 also attains the lowest WAE and WMSE, outperforming both the earlier pipelines and the comparison model C. Not only does this suggest that the design principles in V3 generalize well across different election years, but it also demonstrates that structured reasoning, richer contextual inputs, and ideological self-placement can improve forecasting accuracy and reduce skew even when compared to other leading approaches.
Future Direction 1: Addressing Embedded Political Skewness in Pretrained Corpora. The persistent Democratic skew in simpler pipelines and the residual bias in V3 hint at deeper sources of imbalance in the pretrained corpus. These biases could arise from uneven coverage of political perspectives or disproportionate representation of certain ideologies in the training data Jenny et al. (2024). Addressing this challenge involves examining the corpus composition, adopting balanced data selection strategies, and exploring model-level techniques—such as adversarial debiasing or targeted prompt engineering—to mitigate inherent skew. By addressing these root causes, future election forecasting models can achieve more balanced and trustworthy predictions, ultimately providing a more reliable tool for political analysis and decision-making Li et al. (2024a).
3.2 Result Analysis on LLM-Simulated Voting Patterns
Beyond examining aggregated errors and directional biases, it is important to determine whether the LLM’s predictions capture the demographic patterns observed in real-world voting behavior. We focus on four demographic dimensions—gender, ethnicity, age, and education—identified by Pew Research Center’s 2020 study, Behind Biden’s 2020 Victory Center (2021), as pivotal factors influencing voter preferences. To facilitate this comparison, we aggregated over 200,000 simulated personas across all selected states in 2020, categorizing each persona as either Republican or Democratic based on their predicted votes (excluding non-preference votes). We then examined how these personas’ preferences aligned with the demographic patterns reported by Pew.
(a) Gender and Race Gap: LLM Simulated vs. Pew Report
(b) The Age Gap: LLM Simulated vs. Pew Report
(c) The Education Gap: LLM Simulated vs. Pew Report
Figure 4: LLM-simulated voting gaps across various demographic dimensions: race and gender (a), age (b), and education (c). Each subfigure shows how the LLM’s predictions vary across key demographic factors.
Figure 4 illustrates these comparisons. Directionally, the LLM aligns with Pew’s observations: men, white voters, older adults, and those with lower educational attainment lean more Republican, mirroring established voting patterns. However, the magnitude of these differences is substantially overstated. For instance, while Pew’s data suggest that males only had about a 4% margin in favor of Trump and females about an 11% margin in favor of Biden, the LLM’s simulations inflate these margins to 47.6% among males and 45.6% among females. Similarly, Pew’s data show more modest differences across racial and educational groups, whereas the LLM produces far larger gaps, implying a tendency to reinforce stereotypes rather than reflect real-world proportions.
Future Direction 2: Mitigating the Exacerbation of Stereotypical Biases. While the LLM’s ability to replicate the directional trends found in real-world data is encouraging, the large overemphasis on these demographic distinctions raises concernsChang et al. (2024). Such exaggerations risk reinforcing harmful stereotypes and potentially misrepresenting demographic groups. Moving forward, research should concentrate on developing techniques to calibrate LLM outputs, refine prompt designs, or incorporate counterbalancing information, all aimed at producing simulations that are both directionally accurate and proportionally realisticPark et al. (2024). Ensuring that LLM-based electoral forecasts maintain fairness, rather than amplifying stereotypes, is a critical step toward building more ethical and reliable computational social science tools.
4 Related Work
Applying LLMs in political science represents a new and rapidly evolving field. Although LLMs have revolutionized natural language processing and a variety of other domains, their potential in political science remains largely untapped. Initial studies have demonstrated promising results in election forecasting, policy analysis, and public opinion simulation Smith and Doe (2023); Johnson and Lee (2024). However, political science involves complex social dynamics and multilayered causal relationships, which pose significant challenges to effectively utilize LLMs Brown (2023). Recent work, e.g., Chen and Rodriguez (2024), emphasizes the need for domain-specific fine-tuning to accurately capture the nuances of political discourse, highlighting the limitations of general-purpose LLMs.
LLMs’ applications in election forecasting is still in its early stages. For instance, the “Political Campus” project by Roberts et al. (2023) created a benchmark dataset for evaluating LLM performance on election-related tasks, highlighting both potential and limitations in political forecasting. Similarly, the work by Kim and Patel (2024) using LLMs for policy sentiment analysis underscores the need for careful interpretation of results due to the challenges associated with the inherent complexity of political language. Our research aims to address these challenges by integrating LLMs with ABMs to predict election outcomes and comprehensively analyze voter behavior. This integrated framework combines macro-level political discourse analysis with micro-level voter behavior modeling, bridging the gap identified by previous studies such as Lopez and Singh (2023) regarding the disconnect between large-scale text analysis and individual-level decision-making. For a more comprehensive related work, please see Appx. A.
5 Conclusion and Future Directions
In this work, we present a novel framework for election prediction using large language models (LLMs) with a focus on multi-step reasoning. By leveraging both synthetic personas and real-world datasets, we demonstrated the potential of LLMs to capture individual voting behaviors and state-level election outcomes. Our iterative design highlights the importance of integrating temporal information and complex reasoning for accurate predictions. The 2020 and 2024 simulations reveal both the strengths and limitations of using LLMs in dynamic political environments. These results highlight the model’s ability to generalize on unseen data, while also underscoring several challenges, including: (1) evaluating and addressing biases embedded within the LLM’s pretrained corpus, (2) mitigating stereotypical and overly simplistic representations in persona simulations, and (3) overcoming static demographic assumptions and the lack of real-time data integration.
Future research can extend this work by incorporating multiple LLMs to better understand their internal political tendencies, enhancing temporal modeling with public opinion data and real-time trends for improved accuracy, and developing stronger multi-step reasoning pipelines through refined Chain of Thought (CoT) designs to further enhance prediction performance and mitigate biases. This study lays the foundation for future applications of LLMs in political forecasting, offering promising directions for further development in both election prediction and the broader study of LLM behavior in social contexts.
Broader Impact Statement
This work explores the application of LLMs to improve election forecasting through enhanced reasoning and data synthesis, which can inform policymakers, researchers, and journalists, aiding them in understanding voter behavior and electoral outcomes. By offering a more transparent and adaptable approach to prediction, this research may help demystify complex political processes, reduce reliance on narrow historical data, and guide strategic resource allocation for stakeholders.
Ethics Statement
This work uses only public or synthetic data with no personally identifiable information. Consequently, an institutional review board determined the study to be exempt from further review. While our framework aims to enhance election forecasting accuracy, no model is entirely free of bias. We encourage stakeholders to interpret results carefully, consider domain expertise, and remain vigilant in identifying and mitigating potential biases. Also note that we used ChatGPT exclusively to improve minor grammar in the final manuscript.
Limitations
First, our approach to time-dependent modeling does not account for dynamic factors such as shifts in public opinion, evolving media narratives, or unforeseen events—such as economic disruptions or political scandals—that could significantly influence voter behavior. Incorporating these variables would require real-time data integration, which is beyond the scope of this study.
Second, while our analysis reveals potential political biases within the LLM’s pretrained corpus, we have not quantitatively measured or compared the intrinsic political tendencies of different LLMs. Without a systematic framework to quantify and contrast these inclinations, it remains challenging to understand how model-specific political leanings might influence predictions and outcomes.
It is key to emphasize that this study does not seek to provide definitive election forecasts. Rather, it serves as an exploratory effort to demonstrate the potential of LLMs to generalize to unseen data. By evaluating how closely these models align with real-world patterns, this work offers a proof of concept for their predictive capacity while acknowledging the limitations inherent in the current framework.
Supplementary Material
Appendix A Extended Related Work
A.1 LLMs in Political Science: A New and Emerging Field
The application of LLMs in political science represents a new and rapidly evolving field, with a limited but growing body of research. While LLMs have revolutionized natural language processing and various other domains, their potential in political science remains largely untapped. Initial studies have demonstrated promising results in areas such as election forecasting, policy analysis, and public opinion simulation Smith and Doe (2023); Johnson and Lee (2024). However, political science often involves complex social dynamics and multi-layered causal relationships, posing significant challenges for effectively utilizing LLMs in this context Brown (2023).
Recent work by Chen and Rodriguez (2024) highlights the potential of LLMs in decoding political speeches and policy documents, emphasizing the need for domain-specific fine-tuning to capture the subtleties of political language. Future research is likely to focus on designing models that can handle the intricacies of political discourse while ensuring robustness against biases and misleading inferences Wilson et al. (2024); Thompson (2023).
The latest work by Li et al. (2024a) presents the first systematic framework for integrating LLMs into computational political science. The authors propose a novel taxonomy that categorizes existing work into two main perspectives: political science applications and computational methodologies. From the political science perspective, they highlight LLMs’ capabilities in automating predictive and generative tasks, simulating behavior dynamics, and enhancing causal inference. From the computational perspective, they detail advances in data preparation, fine-tuning approaches, and evaluation methods specifically tailored for political contexts. The paper identifies critical challenges, including the need for domain-specific datasets, addressing bias and fairness issues, incorporating human expertise effectively, and developing evaluation criteria that are aligned with political science requirements. Our work differs from Political-LLM by giving more in-depth analysis on the election forecasting.
A.2 Political Election Research: Classical and Agent-Based Approaches
Traditional political science literature has long relied on survey data and statistical models to analyze voter behavior and predict election outcomes. Classical models like the Downsian spatial model and the Median Voter Theorem explain election results by assuming voters’ positions in policy space Downs (1957); Black (1948). Time-series regression models also play a key role in long-term election forecasting, analyzing the relationships between economic indicators, political events, and voter sentiment Lewis-Beck and Rice (1992); Erikson and Wlezien (2016).
In recent years, Agent-Based Models (ABMs) have gained traction in election research. These bottom-up approaches simulate individual voter decisions and social interactions, providing a more granular view of electoral dynamics. For instance, Gao et al. (2022a) introduced an agent-based election prediction model that captures social networks and voter-candidate interactions to offer more accurate election forecasts. This approach builds upon earlier work by Lemos et al. (2019), who demonstrated the effectiveness of ABMs in modeling voter turnout and preference formation.
ABMs excel at modeling heterogeneity and dynamic processes in real-world voting scenarios, offering greater flexibility and granularity than traditional statistical models. The work of Collins and Martinez (2023) further illustrates how ABMs can incorporate complex factors such as media influence and peer effects in voter decision-making processes. As the field evolves, future election studies may integrate ABMs with LLMs to balance large-scale data analysis with individual voter behavior modeling, as proposed by Zhang et al. (2024b) in their hybrid forecasting framework.
A.3 LLMs in Political Science: Current Research and Our Contributions
Although LLMs have seen rapid advancements, their application in political science remains limited. Only a small number of studies have explored how LLMs can be used for tasks like election prediction, policy analysis, and public opinion tracking. Political text analysis is an important area, and some early benchmark datasets are starting to emerge. However, political language is often complex, with nuanced meanings and context, which presents a significant challenge for LLMs Anderson and Taylor (2023); Williams et al. (2024).
One notable example is the “Political Campus” project by Roberts et al. (2023), which developed a benchmark dataset specifically for election prediction and evaluated LLM performance on various election-related questions. This work has been instrumental in highlighting both the potential and limitations of LLMs in political forecasting. Similarly, the research by Kim and Patel (2024) on using LLMs for policy sentiment analysis demonstrates the models’ capacity to process large volumes of public opinion data, while also underscoring the need for careful interpretation of results.
In contrast to the existing work, our research provides a more comprehensive approach. In 2024, we introduced an integrated framework that combines LLMs with Agent-Based Models to predict election outcomes and analyze voter behavior. Unlike previous studies, our approach not only focuses on semantic analysis of political texts but also incorporates individual voter behavior modeling, offering more granular and accurate election predictions. This integration addresses the limitations identified by Lopez and Singh (2023) regarding the disconnect between macro-level language models and micro-level voter behavior.
Appendix B Details of Multi-Step LLM Pipeline
B.1 Dataset Details
B.1.1 Real-world Data by American National Election Studies (ANES)
For evaluation, we use pre-election data from the ANES 2016 and 2020 Time Series Studies Argyle et al. (2022); Studies (2022), which provide 4,270 and 8,280 real-world samples, respectively, from individuals who participated in the 2016 and 2020 elections. The dataset includes a wide range of variables: (1) racial/ethnic self-identification, (2) gender, (3) age, (4) ideological self-placement on a conservative-liberal scale, (5) party identification, (6) political interest, (7) church attendance, (8) frequency of discussing politics with family and friends, (9) patriotic feelings associated with the American flag (unavailable in 2020), and (10) state of residence (unavailable in 2020). Additionally, the dataset records how individuals voted in both the 2016 and 2020 elections. Previous studies, such as Argyle et al. (2023), have evaluated GPT-3 using this dataset. We apply our method directly to this established benchmark to assess its effectiveness and performance.
B.1.2 Synthetic Data for the US Population
In addition to the medium-sized benchmark dataset, we utilize synthetic demographic data derived from a 1:1 synthetic population dataset of the United States Li et al. (2020b). Synthetic data plays a crucial role in social and applied sciences, with recent applications in water quality estimation Chia et al. (2023), financial modeling Potluru et al. (2023a), tourist profiling Merinov et al. (2023a), and measuring the social impact of engineered products Stevenson et al. (2023). High-quality synthetic datasets provide researchers with large-scale data at a lower cost while maintaining privacy, making them a reliable resource.
For our purposes, the synthetic data enables the creation of a cost-effective, large-scale virtual panel of respondents that is both “wide" (each respondent has over 50k modeled features) and “long" (enough samples to reflect a national dataset). However, running LLM inference on the entire U.S. population would be prohibitively expensive, so we employ a sampling strategy. Given the pivotal role of swing states in determining election outcomes, we focus on simulating voter behavior in these states while including representative samples from red and blue states for comparison.
Synthetic Data Generation: The synthetic data used here is generated using the SynC framework Li et al. (2020b), which reconstructs individual-level data from aggregated sources where collecting real-world individual data is impractical due to privacy, time, or financial constraints. SynC is widely recognized and applied across multiple fields to support research and overcome data limitations. For instance, it has been used in outlier detection Li et al. (2020a), finance Potluru et al. (2023b), tabular data modeling Borisov et al. (2022), healthcare Sichani et al. (2024), and tourism Merinov et al. (2023b), demonstrating its effectiveness and importance in various domains.
SynC leverages publicly available data, such as the 2023 American Community Survey (ACS), which provides data on 242,338 census block groups, including population statistics and response proportions for each block. Using Data Downscaling, SynC probabilistically recreates the 340 million residents represented in the aggregated census data. For our simulation, the synthetic population includes variables relevant to election predictions: (1) age, (2) gender, (3) ethnicity, (4) marital status, (5) household size, (6) presence of children, (7) education level, (8) occupation, (9) individual income, (10) family income, and (11) place of residence.
SynC addresses the challenge of reconstructing individual data $\{x_{m,1}^{d},\ldots,x_{m,n_{m}}^{d}\}$ from aggregated observations $X_{m}^{d}=\sum_{k=1}^{n_{m}}x_{m,k}^{d}/n_{m}$, where $X^{d}$ is the $d$-th survey question of interest, $m$ is the census block id and $n$ is the number of individuals in $m$. A Gaussian copula is employed to model dependencies between survey questions. Given a $d\times d$ covariance matrix $\Sigma$ of the $d$ sruvey questions, the synthetic individuals are drawn as:
$Z_{m}^{d}\sim N(0,\Sigma),\quad u_{m}^{d}=\Phi(Z_{m}^{d}),\quad X_{m}^{d}=F^{-1}_{d}(u_{m}^{d}),$ (A1)
where $Z_{m}^{d}\sim N(0,\Sigma)$ denotes a random seed from a multivariate normal distribution, $\Phi$ is the cumulative distribution function (CDF) of the standard normal distribution, and $F^{-1}_{d}$ is the inverse CDF of the marginal distribution for feature $d$, which is estimated based on census block level data. To maintain alignment with aggregated data, SynC uses marginal scaling. For categorical variables, it applies a multinomial distribution:
$X^{d}\sim\text{Multi}(1,c^{d},p_{m,k}^{d}),$ (A2)
where $p_{m,k}^{d}$ is the probability distribution over $c^{d}$ categories for question $d$ and individual $k$. Marginal constraints are adjusted iteratively if discrepancies arise between sampled and target proportions.
The multi-phase SynC framework ensures that: (1) marginal distributions of individual features align with real-world expectations, (2) feature correlations are consistent with aggregated data, and (3) aggregated results match the input data. For further details on SynC’s methodology and algorithms, please see the original paper Li et al. (2020b).
Partition Design and State Categorization: The synthetic dataset evaluation will operate at the state level, where we sample synthetic individuals from each state to simulate voter behavior and aggregate their votes to compare the simulated outcomes with actual election results. Given the critical role of swing states and tipping-point states in determining election outcomes, our primary focus is on these states, which include Florida (FL), Wisconsin (WI), Michigan (MI), Nevada (NV), North Carolina (NC), Pennsylvania (PA), Georgia (GA), Texas (TX), Minnesota (MN), Arizona (AZ), and New Hampshire (NH). For broader comparison in the following evaluations, we also sample from several reliably “red states,” such as Alabama (AL), Arkansas (AR), Idaho (ID), Ohio (OH), and South Carolina (SC), as well as from “blue states,” such as California (CA), Illinois (IL), New York (NY), New Jersey (NJ), and Washington (WA). These classifications are based on the 2020 election results as described by Wikipedia contributors (2024).
Sampling Method: Running LLM inference on the entire synthetic population is computationally prohibitive, so we adopt a random sampling approach. Each state serves as a sampling unit, with sample sizes ranging between 1/100 and 1/2000 of the synthetic population, depending on the state’s population size. For example, a 1/2000 sampling ratio is applied to highly populated states like California, while a 1/100 ratio is used for smaller states such as New Hampshire. This approach ensures a minimum sample size of $4269$ individuals per state, corresponding to a 1.5% margin of error at a 95% confidence level, to maintain sufficient representation. Although our primary focus is on swing states due to their critical influence on election outcomes, we apply the same sampling method to red and blue states included in our simulations to ensure consistency across the analysis.
B.2 Detailed Evaluation Metrics
To comprehensively evaluate our proposed approaches, we employ multiple metrics for both benchmark datasets (ANES 2016 and 2020 Studies (2019, 2022)) and state-level simulations. For the ANES benchmarks, we follow the methodology of Argyle et al. (2023), comparing the average voting probabilities:
1. Predicted Proportion ($P(s)$)
Probability=Republican VotesRepublican Votes+Democratic $\text{Probability}=\frac{\text{Republican Votes}}{\text{Republican Votes}+\text{Democratic Votes}}$ (A3)
For state-level comparisons, we introduce the following additional metrics:
2. Weighted Absolute Error (WAE)
$\text{WAE}=\frac{\sum_{s\in S}E(s)\cdot|P(s)-R(s)|}{\sum_{s\in S}E(s)}$ (A4)
where:
  • $P(s)$: The predicted proportion, calculated as the ratio of Republican votes to total votes (Republican + Democrat) for each state (A3).
  • $R(s)$: The actual proportion of votes in state $s$.
  • $E(s)$: The electoral votes assigned to state $s$, serving as weights.
  • $S$: The set of all selected states.
3. Weighted Mean Squared Error (WMSE)
$\text{WMSE}=\frac{\sum_{s\in S}E(s)\cdot(P(s)-R(s))^{2}}{\sum_{s\in S}E(s)}$ (A5)
where:
  • $(P(s)-R(s))^{2}$: The squared error between the predicted and actual proportions for each state.
4. Bias Metric (BM)
$\text{BM}=\frac{\sum_{s\in S}E(s)\cdot(P(s)-R(s))}{\sum_{s\in S}E(s)}$ (A6)
where:
  • Positive Value: Reflects a systematic overestimation of $P(s)$, indicating a bias toward the Republican Party.
  • Negative Value: Reflects a systematic underestimation of $P(s)$, indicating a bias toward the Democratic Party.
These metrics are calculated across the entire sample to evaluate both the magnitude and direction of errors. Accuracy is further assessed by comparing the predicted winning party with the actual election outcome.
For the synthetic dataset, we treat each state as an independent validation unit. The predicted results—both in terms of the winning candidate and vote share percentages—are compared against the actual 2020 election results for each state. Accuracy is evaluated based on:
  1. 1.
    Agreement between the predicted and actual winning candidate for each state.
  2. 2.
    Aggregate performance across all states, ensuring the model captures overall election trends.
This state-level evaluation leverages voter-level information processed through LLMs to generate accurate predictions, providing a robust assessment of model performance across diverse electoral scenarios.
Appendix C Evaluations on Synthetic Data for the 2020 US Population
Figure A1: LLM’s predictions for four states in the 2020 election compared with Ground Truth results. The figure presents results for one red state (Ohio, OH), one blue state (Illinois, IL), one swing state (Wisconsin, WI), and one tipping-point state (Florida, FL). V1 and V2 pipelines tend to underestimate Republican support, while V3 (Multi-step Reasoning) provides the closest alignment with actual outcomes, especially in swing and tipping-point states.
In addition to the nationwide evaluation on the ANES datasets, we conducted state-level simulations using synthetic data to compare predictions with actual 2020 election outcomes. For each state, we performed random sampling based on population size to ensure a statistically meaningful number of personas. The simulation outcomes were then benchmarked against official 2020 Presidential General Election Results from the Federal Election Commission (FEC). As in the benchmark evaluations, we calculated the average voting probabilities to assess the alignment of predictions with real-world outcomes. We evaluated five red states, five blue states, and 11 swing and tipping-point states. Figure A1 highlights representative results from these categories, providing insights into the model’s performance in different electoral contexts.
Figure A2: Aggregated results of the three pipelines (V1, V2, V3) on state-level simulations. Each confusion matrix presents the number of states where predictions align with or deviate from actual outcomes. V1 (AUC = 0.69) and V2 (AUC = 0.62) show lower accuracy, while V3 (AUC = 0.90) performs best, effectively capturing Republican victories without compromising Democratic predictions. It is worth noting that, so far, we have only tested the pipelines in 21 states. If the scope is expanded to include all states, the AUC of V3 is expected to improve further, while the AUC of V1 and V2 are expected to decline.
Consistent with the ANES dataset evaluations, the V1 pipeline (Demographic-only Prompt) exhibited a skew toward the Democratic Party, even in traditionally Republican-leaning states like South Carolina (SC), Alabama (AL), and Ohio (OH), with predictions diverging significantly from actual results. This illustrates the limitations of using demographic data alone without time-sensitive context. The V2 pipeline (Time-dependent Prompt) introduced election-year-specific information, which partially reduced the skew in the state-level simulations. However, the model still struggled to eliminate prediction biases, particularly in polarized states. Interestingly, this differed from the ANES evaluations, where including time-dependent information amplified the bias. The V3 pipeline (Multi-step Reasoning) demonstrated the most accurate performance, effectively mitigating skewness across deep red and blue states. In these polarized states, the predictions closely mirrored the actual voting outcomes, reflecting the model’s improved ability to incorporate ideological alignment through multi-step reasoning.
For swing and tipping-point states, the V3 pipeline achieved robust results, correctly predicting the outcomes in 9 out of 11 states. Minor deviations were observed in North Carolina (NC) and Arizona (AZ), where the predictions were slightly misaligned with the real results. Nonetheless, the V3 pipeline provided balanced predictions that accurately captured the competitive dynamics typical of swing states, further validating its effectiveness.
In summary, the comparative performance of the three pipelines across different state categories is shown in Figure A1. The V3 pipeline consistently outperformed the other two, delivering more stable and accurate predictions. Aggregate results for all pipelines on all 21 chosen states is shown in the below figure  A2. And all state-level simulation results can be found in the Figure A3.
Appendix D Additional Results on 2024 Prediction
The 2024 state-level predictions offer deeper insights into the performance of the proposed pipelines across diverse electoral contexts. As discussed in §3.1, simulations for the 2024 election indicate a systematic bias toward the Democratic Party across the 11 swing and tipping-point states. This bias may reflect the LLM’s sensitivity to candidate-specific factors in the 2024 context. Specifically, prior to the election, V3 (§2.2.3) was evaluated across all 50 states, while V1 (§2.2.1) and V2 (§2.2.2) were tested on selected swing states and traditional red and blue states. The comparative performance of these pipelines is presented in Figure A4.
At the state level, several notable shifts are observed compared to 2020 predictions. For instance, Wisconsin (WI) demonstrates a significant change, with Trump projected to win 54.90% of the vote. Gains are also observed in Pennsylvania (PA) 47.85%, Michigan (MI) 48.87%, and New Hampshire (NH) 49.49%, though these states remain highly competitive.
In other key battleground states, Arizona (AZ) is forecasted to return to the Republicans with 51.09%, while Florida (FL) and Texas (TX) continue to show strong Republican support at 53.62% and 56.36%, respectively. Conversely, in contrast to the actual results, Nevada (NV) at 34.77%, Georgia (GA) at 44.36%, and Minnesota (MN) at 42.95% are predicted to lean more toward the Democratic Party, highlighting the complex dynamics of these closely contested regions.
In traditional strongholds, the predictions align with historical trends. Republican-dominated states like Arkansas (AR) and Alabama (AL) continue to show robust GOP support, while Democratic bastions such as California (CA), New York (NY), and Illinois (IL) remain reliably blue. An exception is Alaska (AK) 49.39%, where the model predicts a closer contest compared to prior elections.
These state-level results, summarized in Figure A4, highlight the nuanced performance of the pipelines. The varying prediction patterns underscore both the strengths and limitations of the models, emphasizing opportunities for further refinement to better capture the complexities of voter behavior and electoral dynamics.
Figure A3: Overall performance of the three pipelines (V1, V2, V3) in the 2020 election across all selected states. These include five red states (Arkansas (AR), South Carolina (SC), Idaho (ID), Alabama (AL), Ohio (OH)), five blue states (Illinois (IL), California (CA), New York (NY), New Jersey (NJ), Washington (WA)), 11 swing and tipping-point states (New Hampshire (NH), Arizona (AZ), Pennsylvania (PA), Georgia (GA), Minnesota (MN), North Carolina (NC), Wisconsin (WI), Florida (FL), Michigan (MI), Nevada (NV), Texas (TX)), and an additional red state (Alaska (AK)). The red reference line corresponds to the official 2020 election results published by the Federal Election Commission Federal Election Commission (2021), while the black reference line represents an equal vote share (0.5) between the two parties.
Figure A4: Overall performance of the three pipelines (V1, V2, V3) in the 2024 election across five red states (Arkansas (AR), South Carolina (SC), Idaho (ID), Alabama (AL), Ohio (OH)), five blue states (Illinois (IL), California (CA), New York (NY), New Jersey (NJ), Washington (WA)), 11 swing and tipping-point states (New Hampshire (NH), Arizona (AZ), Pennsylvania (PA), Georgia (GA), Minnesota (MN), North Carolina (NC), Wisconsin (WI), Florida (FL), Michigan (MI), Nevada (NV), Texas (TX)), and an additional red state (Alaska (AK)). The red reference line corresponds to the 2024 election results reported by the NBC News NBC News (2024), while the black reference line represents an equal vote share (0.5) between the two parties.

### v3

@@
-A Large-scale Empirical Study on Large Language Models for Election Prediction
+A Large-Scale Simulation on Large Language Models for Decision-Making in Political Science
 Abstract
-Can Large Language Models (LLMs) accurately predict election outcomes?
-While LLMs have demonstrated impressive performance in healthcare, legal analysis, and creative applications, their capabilities in election forecasting remain uncertain.
-Notably, election prediction poses unique challenges: limited voter-level data, evolving political contexts, and the complexity of modeling human behavior.
-In the first part of this paper, we explore and introduce a multi-step reasoning framework for election prediction, which systematically integrates demographic, ideological, and time-sensitive factors.
-Validated on 2016 and 2020 real-world data and extensive synthetic personas, our approach adapts to changing political landscapes, reducing bias and significantly improving predictive accuracy.
-We further apply our pipeline to the 2024 U.S.
-presidential election, illustrating its ability to generalize beyond observed historical data.
-Beyond enhancing accuracy, the second part of the paper provides insights into the broader implications of LLM-based election forecasting.
-We identify potential political biases embedded in pretrained corpora, examine how demographic patterns can become exaggerated, and suggest strategies for mitigating these issues.
-Together, this project, a large-scale LLM empirical study, advances the accuracy of election predictions and establishes directions for more balanced, transparent, and context-aware modeling in political science research and practice.
-A Large-scale Empirical Study on Large Language Models for Election Prediction
-Chenxiao Yu1, Zhaotian Weng1, Yuangang Li1, Zheng Li2, Xiyang Hu3, Yue Zhao1 1University of Southern California  2Arima  3Carnegie Mellon University {cyu96374, wengzhao, yuangang, yzhao010}@usc.edu, winston@arimadata.com, xiyanghu@cmu.edu Corresponding authors
+While LLMs have demonstrated remarkable capabilities in text generation and reasoning, their ability to simulate human decision-making—particularly in political contexts—remains an open question.
+However, modeling voter behavior presents unique challenges due to limited voter-level data, evolving political landscapes, and the complexity of human reasoning.
+In this study, we develop a theory-driven, multi-step reasoning framework that integrates demographic, temporal and ideological factors to simulate voter decision-making at scale.
+Using synthetic personas calibrated to real-world voter data, we conduct large-scale simulations of recent U.S.
+presidential elections.
+Our method significantly improves simulation accuracy while mitigating model biases.
+We examine its robustness by comparing performance across different LLMs.
+We further investigate the challenges and constraints that arise from LLM-based political simulations.
+Our work provides both a scalable framework for modeling political decision-making behavior and insights into the promise and limitations of using LLMs in political science research.
+A Large-Scale Simulation on Large Language Models for Decision-Making in Political Science
+Chenxiao Yu1, Jinyi Ye1, Yuangang Li1, Zhaotian Weng1, Zheng Li2, Emilio Ferrara1, Xiyang Hu3, Yue Zhao1 1University of Southern California  2Arima  3Arizona State University {cyu96374, jinyiy, yuangang, wengzhao, emiliofe, yzhao010}@usc.edu winston@arimadata.com, xiyanghu@asu.edu Corresponding authors
 1 Introduction
-Large Language Models (LLMs) have shown strong capabilities across diverse domains, including drug discovery Liu et al.
-(2024), legal analysis Chalkidis et al.
-(2022), and creative tasks Yang et al.
+Large Language Models (LLMs) have demonstrated strong capabilities in processing and generating text, drawing on vast amounts of knowledge to assist with tasks in fields like scientific discovery Liu et al.
+(2024), law Chalkidis et al.
+(2022), and creative work Yang et al.
 (2022).
-Their strength lies in not only understanding and generating text but also leveraging vast common knowledge Roberts et al.
-(2020), simulating diverse personas Hu and Collier (2024), and modeling human behavior in complex social contexts Bommasani et al.
+Beyond text generation, they show emerging reasoning abilities that allow them to approximate human-like thought processes Zhou et al.
+(2020); AlKhamissi et al.
+(2022) and model human behavior Bommasani et al.
 (2021).
-LLMs excel in human-like reasoning Zhou et al.
-(2020); AlKhamissi et al.
-(2022) and have been applied to simulate decision-making in various settings Zhou et al.
-(2023); Ziems et al.
+However, LLMs still struggle to capture the deeper psychological and social mechanisms that drive human decision-making, making their simulations less reliable in real-world contexts Zhou et al.
 (2024).
-Notably, recent work has extended LLMs to political science, analyzing policy documents, campaign speeches, and public sentiment Xu (2022); Haq et al.
-(2023); Li et al.
-(2024b).
-Beyond language processing, their ability to integrate knowledge and apply human-like reasoning positions them as powerful tools for simulating the intricate dynamics of political decision-making Argyle et al.
-(2023); Bisbee et al.
+To address this, researchers have started incorporating insights from social science into LLM-based models.
+Recent studies have explored how LLMs can simulate economic decision-making Ross et al.
+(2024), public opinion dynamics Chuang et al.
+(2024), and social-psychological mechanisms like collaboration and conformity Zhang et al.
+(2024a).
+Despite these advances, the application of LLMs to political decision-making remains underexplored.
+Voting behavior is one of the most fundamental decision-making processes in political science.
+LLMs are well-suited for this task because they have shown strong zero- and few-shot capabilities in simulating human dynamics, like political homophily in social networks Chang et al.
 (2024).
-Figure 1: Demonstration of three prompt designs in §2.2.
-V1 is the direct prompt on voter demographic information, while V2 introduces time-dependent information to capture candidates’ agenda and V3 also uses multi-step reasoning.
-In this example for 2020 Ohio result prediction, only V3 can accurately predict the results, demonstrating the importance of leveraging both time-dependent information and multi-step reasoning for election result prediction.
-Background.
-Despite LLMs’ achievements in more straightforward political science tasks, their capacity to handle more challenging efforts, such as election prediction, is not well understood Lerer et al.
+Just as they have been used to estimate politicians’ ideological positions Wu et al.
+(2023), they could also approximate the average voter’s behavior, providing a scalable way to analyze political preferences at a broader level.
+Election simulations naturally emerge as a structured application of this approach.
+Unlike abstract ideological simulations, election outcomes offer a clear ground truth—real-world state- and county-level election data—making election simulation an ideal testbed for evaluating LLMs’ reasoning and predictive abilities in political science.
+If successful, this approach could extend to downstream applications, such as forecasting public reactions to policy changes, where traditional large-scale surveys and experiments are often costly and time-consuming.
+Yet, accurately simulating voting behavior presents a number of challenges.
+First, LLMs inherit political biases from the data they are trained on, which can skew their predictions in politically sensitive tasks Feng et al.
 (2023).
-LLMs’ access to extensive historical data suggests potential for forecasting, but election prediction introduces unique challenges.
-First, the high cost of acquiring voter-level data complicates both experimentation and model validation.
-Second, election forecasting involves modeling not only voter behavior, but also the shifting political context, making it unclear whether text-based data alone can capture the full complexity Graefe (2014).
-Third, accurate predictions require reasoning about multiple factorsHolbrook (2016).
-Whether LLMs can perform this more advanced reasoning remains unknown.
-Wei et al.
+Second, voting decision-making is shaped by various factors, including demographics, location, ideology, and party affiliation Levendusky (2009); Abramowitz and Saunders (2008), but the high cost of acquiring voter-level data complicates both experimentation and model validation.
+Third, a large-scale election simulation should account not only for individual voter behavior but also for the shifting political context, but it remains unclear whether text-based data alone is sufficient to capture these information Graefe (2014).
+Fourth, accurate simulations may require multi-step reasoning Holbrook (2016), yet how to effectively integrate political science insights into LLMs’ reasoning processes—and whether LLMs can handle this level of complexity—remains an open question Wei et al.
 (2022).
 This Work.
-We present a large-scale empirical study to address two key questions in election prediction with LLMs: (i) how can LLMs be leveraged to achieve accurate election predictions, and (ii) beyond accuracy, what additional considerations are crucial for understanding LLM capabilities in election forecasting tasks?
-Contribution 1: Multi-Step Reasoning Framework for Accurate Election Prediction (§2).
-To handle the scarcity of detailed voter-level data, we employ the Sync synthetic data generation framework Li et al.
+We present a large-scale simulation study exploring how LLMs can model human decision-making in political science, focusing on voter behavior in U.S.
+elections.
+We develop a theory-driven, multi-step reasoning framework that incorporates demographic, ideological, and temporal factors to model political decision-making at scale.
+We evaluate our framework on different LLMs, compare their robustness and predictive performance, and investigate biases and limitations that emerge in large-scale political simulations.
+Our study addresses three key research questions:
+RQ1: How can LLMs be used to simulate human decision-making in political science?
+RQ2: How do different LLMs perform, and how robust are their election simulations?
+RQ3: What limitations arise when using LLMs to model political decision-making?
+Contribution 1: A Theory-Driven Multi-Step Reasoning Pipeline for Accurate Election Simulation (§2).
+We propose a theory-driven, multi-step reasoning pipeline to simulate voter decision-making, incorporating demographic, temporal, and ideological factors.
+To address the lack of detailed voter-level data, we use the Sync synthetic data generation framework Li et al.
 (2020b), which probabilistically reconstructs individual demographic and behavioral profiles from aggregated public datasets.
-We pair this synthetic data with real-world datasets, notably the American National Election Studies (ANES) Time Series Studies (2022), ensuring that our model aligns with actual voting behaviors.
-Our approach also adapts to evolving political conditions by incorporating time-dependent factors.
-We aggregate information from presidential campaign data, including candidates’ policy agendas and backgrounds, to align the model with changing political environments Holbrook (2016).
-Finally, we introduce a multi-step reasoning framework guided by Chain of Thought prompting Wei et al.
+We then align the personas with real-world voter data from American National Election Studies (ANES)  [^1] https://electionstudies.org/.
+Our approach also adapts to evolving political conditions by integrating temporal factors, such as candidates’ policy agendas and backgrounds Holbrook (2016).
+Furthermore, building on political science studies on ideological sorting—the process by which voters increasingly align their political ideology with their party affiliation over time Levendusky (2009), we introduce ideology inference as an intermediate reasoning step.
+Using Chain-of-Thought prompting Wei et al.
+(2022), our model first predicts ideology based on demographics and behavioral data, which then influences party affiliation and voting preferences.
+As shown in Fig 1, we refine our pipeline iteratively, incorporating demographics, political context, and ideological inference at each step.
+The final model significantly improves in simulation accuracy and alignment with real-world election.
+Contribution 2: Challenges and Limitations in Large-Scale Political Simulations (§3).
+Our analysis reveals three important challenges in LLM-based political simulations.
+First, LLMs inherit systematic biases from their pretraining data, leading to a persistent left-leaning skew in simpler pipelines, with multi-step reasoning reducing but not eliminating this bias.
+Second, LLMs exaggerate demographic voting patterns, amplifying stereotypes related to gender, race, and education.
+Third, LLMs overestimate the influence of ideology on voting behavior, producing higher-than-real-world correlations between ideology and voting preference, a phenomenon referred to in previous work as “hyper-accuracy distortion” Aher et al.
 (2023).
-By breaking down the prediction process into intermediate steps, the model can systematically integrate demographic details, ideological alignment, and time-sensitive information.
-This approach strengthens the model’s accuracy and mitigates biases and overfitting issues observed in simpler models.
-As shown in Figs.
-1 and 2, we refine our pipeline through iterations.
-The final version significantly improves predictive accuracy and alignment with real-world data via extensive experiments.
-Contribution 2: New Insights and Directions Beyond Accurate Prediction (§3).
-Our experiments uncover several key insights that extend beyond raw predictive accuracy.
-First, while our multi-step reasoning pipeline reduces bias and improves alignment with real-world outcomes, residual skewness persists, highlighting the influence of pretrained corpora on political predictions.
-Second, although LLMs can replicate known demographic voting patterns, they tend to exaggerate these differences, raising concerns about reinforcing stereotypes.
-Finally, our results indicate that integrating richer contextual information and refining prompt strategies can further enhance predictive reliability.
-We thus propose meaningful future directions on reducing political skewness in pretrained corpora and mitigating stereotype biases in election prediction.
-2 Part 1: Accurate Election Prediction via a Multi-Step LLM Pipeline
-How can we effectively leverage LLMs to predict election results?
+By these findings, we propose future research directions focused on debiasing training data, refining demographic calibration, and introducing human-in-the-loop techniques to improve the accuracy and reliability of LLM-based political simulations.
+2 Accurate Simulation of Human Voting Behavior via a Multi-Step LLM Pipeline (RQ1, 2)
+Figure 1: Progressive design of LLM pipelines for voter simulation.
+V1: Demographic-Only Prompting2.2.1) uses static personas but lacks temporal context.
+V2: Time-Based Prompting2.2.2) adds election-year data V3: Multi-Step Reasoning2.2.3) structures decision-making into steps, improving reasoning and alignment.
+How can we use LLMs to simulate human voting behavior in political science?
 In this work, we simulate each voter’s decision-making process by providing LLMs with detailed voter information and asking them to predict voter preferences.
-To achieve this, we focus on two key aspects: (1) establishing an evaluation framework with appropriate datasets that contain voter-level information, and (2) designing an LLM-based pipeline for accurate election predictions.
-In §2.1, we describe the datasets and the evaluation methods.
-We then provide an overview of our design approach in §2.2 with three progressive pipelines.
-Finally, we evaluate the performance of the pipelines on datasets in §2.3.
-Figure 2: Progressive design of LLM pipelines for election predictions.
-V1: Demographic-only Prompting2.2.1) uses static demographic personas but lacks temporal context.
-V2: Time-based Prompting2.2.2) adds election-year policy shifts and candidate information, but the single-step prompt setup leads to overloaded inputs that reduce prediction accuracy.
-V3: Multi-step Reasoning2.2.3) divides the decision-making process into multiple steps, improving reasoning quality and producing results that closely match real-world outcomes.
-All versions aggregate individual predictions into state-level results to capture broader election trends.
+To achieve this, we focus on two key components: (1) building a robust evaluation framework using datasets that contain voter-level information, and (2) designing a theory-driven Bafumi and Shapiro (2009); Pew Research Center (2014) LLM-based pipeline for accurate election simulation.
+In §2.1, we present the datasets and evaluation methods.
+Next, we outline our design approach in §2.2, introducing three progressive pipelines, incorporating demographics, political context, and ideological inference at each step.
+Finally, we evaluate these pipelines by comparing their predictions with real-world outcomes in §2.3.
 2.1 Datasets, Evaluation, and Settings
 Datasets.
-This study leverages two primary data sources: (1) Public Benchmarks: The American National Election Studies (ANES) 2016 and 2020 Time Series data Studies (2019, 2022), which provide comprehensive demographic and political ideology information for real respondents.
-This dataset is a benchmark to evaluate how accurately LLMs simulate voter-level behavior in alignment with real-world data.
-(2) For state-level voting simulations, we use voter-level synthetic data generated via advanced ML techniques based on aggregated information Li et al.
+This study leverages two primary data sources: (1) Public Benchmarks: The American National Election Studies (ANES) 2016 and 2020 Time Series data Studies (2019, 2022), which provide detailed demographic, ideology, and party affiliation information from real respondents.
+This dataset serves as a benchmark to evaluate how well LLMs simulate voter-level behavior in alignment with real-world patterns.
+(2) Large-Scale State-level Synthetic Voter Persona Dataset: A dataset of over 330,000 synthetic personas, generated using advanced ML techniques based on aggregated population census data and commercial datasets Li et al.
 (2020b).
-Personas are randomly sampled for each state at specified ratios, and predictions are compared with actual election outcomes from 2020 Federal Election Commission (2021) and 2024 NBC News (2024).
-Both datasets contain non-personally identifiable [^1] This project has been reviewed by the IRB and considered exempt as the datasets are non-personally identifiable., voter-level information.
-Detailed partition and sampling methodologies are in Appx.
-B.1.
+Personas are randomly sampled for each state at specified ratios, and their predicted voting outcomes are compared to actual U.S.
+election results from 2020 Federal Election Commission (2021) and 2024 NBC News (2024).
+Both datasets contain non-personally identifiable voter-level information [^2] This project has been reviewed by the IRB and exempt, as the datasets do not include personally identifiable info..
+Detailed partitioning and sampling methodologies are provided in Appx.
+A.1.
 Evaluation Method.
-To evaluate performance on public benchmarks and state-level simulations, we firstly assess how closely LLM simulations align with actual voting results.
-The calculation follows the approach outlined in Xie et al.
-(2024); Argyle et al.
-(2023) and is defined as:
-Predicted Proportion Ratio $P(s)$
-$P(s)=\frac{\text{Republican Votes}}{\text{Republican Votes}+\text{Democratic Votes}}$ (1)
-Here, $s$ represents the unit of analysis: it can refer to an individual persona (as in public benchmarks like ANES) or an entire state (as in state-level simulations).
-This ratio $P(s)$ measures the proportion of votes predicted for the Republican Party relative to the total votes for the two major parties, excluding those with no preference.
-Hardware and LLM Settings.
-Our experiments were conducted on a GPU server with an AMD EPYC Milan 7763 processor with 6 NVIDIA RTX A6000 Ada GPUs.
-For the LLM, we use OpenAI’s GPT-4o model for election predictions.
-Meta’s LLaMA 3.1 (405B) model is employed in intermediate steps to provide neutral summarizations of time-dependent information, enhancing certain pipelines for capturing temporal dynamicsFeng et al.
+To evaluate performance on public benchmarks and state-level simulations, we assess how closely LLM simulations align with actual voting results.
+The calculation follows the approach outlined in Argyle et al.
+(2023):
+Predicted Voting Ratio $P(s)$
+$P(s)=\frac{\textit{Republican Votes}}{\textit{Republican Votes}+\textit{Democratic Votes}}$ (1)
+Here, $s$ represents the unit of analysis, which can refer to cross-regional samples (e.g., public benchmarks like ANES) or an entire state (e.g., state-level simulations).
+The ratio $P(s)$ measures the number of votes predicted for the Republican Party relative to the total votes for the two major parties, excluding those who express no preference.
+LLMs and Hardware Settings.
+Our experiments utilized OpenAI’s GPT-4o and Meta’s LLaMA 3.1-70B model for the primary simulations.
+Meta’s LLaMA 3.1 (405B) model was employed in intermediate steps to provide neutral summarizations of time-dependent information Feng et al.
 (2023).
+Furthermore, we tested Qwen-72B and DeepSeek-V3 to measure systematic differences between models.
+For the hardware setup, we employed six NVIDIA RTX A6000 Ada GPUs and an 8-way NVIDIA A100 GPU cluster, with AMD Milan processors to execute tasks across different models.
 2.2 Our Progressive Design of LLM Pipelines
-In this section, we present our progressive design for generating voter-level election predictions using LLMs.
+In this section, we present our progressive design for generating voter-level behavior simulation using LLMs.
 As shown in Fig.
-2, we develop three versions of the pipeline.
+1, we develop three versions of the pipeline.
 Each version addresses a key shortcoming of its predecessor and integrates more detailed information and reasoning processes.
   1. V1:
-    Demographic-only Prompting (§2.2.1): This initial version uses static demographic personas to guide voter-level predictions.
-It is simple but does not capture changes in candidates’ focus over time.
+    Demographic-Only Prompting (§2.2.1): This baseline approach uses static demographic personas for voter-level simulations.
+While straightforward, it does not account for shifts in presidential candidates’ policy priorities over time.
   2. V2:
-    Single-step Prompting with Time-based Information (§2.2.2): Here, we add election-year-specific details, such as policy agendas and candidate backgrounds, to make predictions more dynamic.
-However, packing all information into a single prompt can overwhelm the model, limiting its reasoning depth.
+    Single-Step Prompting with Time-Sensitive Information (§2.2.2): Here, we add election-year-specific details, like policy agendas and candidate backgrounds.
+However, packing all information into a single prompt may overwhelm the model, limiting reasoning depth.
   3. V3:
-    Multi-step Reasoning with Contextual Information (§2.2.3): This version separates the prediction process into a sequence of steps.
-By breaking down the reasoning, the model can integrate demographics, candidate profiles, and political context better, for predictions that better match real-world results.
-2.2.1 Version 1: Demographic-only Prompting
-This initial version directly prompts the LLM with a persona that includes demographic attributes (e.g., age, gender, income) and asks how that persona would vote Xie et al.
-(2024); Argyle et al.
+    Multi-Step Reasoning with Ideology Inference (§2.2.3): This version structures the simulation into sequential steps, allowing the model to better integrate demographics, political ideology, and political context for more accurate real-world predictions.
+2.2.1 V1: Demographic-Only Prompting
+This initial version prompts the LLM with a persona’s demographics (e.g., age, gender, income) to simulate voting behavior Argyle et al.
 (2023).
+To prevent confusion, we specify the year as 2020, ensuring alignment with the LLM’s training data, which extends through 2023.
+The listed voting options follow Pew Research Center’s 2014 Political Polarization and Typology Survey Pew Research Center (2014).
 Task: You are persona [age, gender, ethnicity, marital status, household size, presence of children, education level, occupation, individual income, family income, and place of residence.] The current year is [year].
-Please answer the following question as if you were the resident: 1. As of today, will you vote for the Democratic Party (Joe Biden), the Republican Party (Donald Trump), or do you have no preference?
+Please answer the following question as if you were the resident:
+1. As of today, will you vote for the Democratic Party (Joe Biden), the Republican Party (Donald Trump), or do you have no preference?
 Options: Democratic, Republican, No Preference
-Specifying the year as 2020 avoids confusion because the LLM’s knowledge ends in 2023.
-The listed voting options follow Pew Research Center’s 2014 Political Polarization and Typology Survey Pew Research Center (2014).
-Limitations: This version cannot adapt predictions to different time periods.
-Without incorporating temporal changes in candidates’ agendas or public opinion, predictions remain static, limiting their usefulness for evolving electoral contexts.
-2.2.2 Version 2: Single-step Prompting with Time-based Information
+Limitations: This version lacks adaptability to different election cycles.
+Without accounting for shifts in candidate agendas or public opinion, its predictions remain static, limiting relevance in changing electoral contexts.
+2.2.2 V2: Single-Step Prompting with Time-Sensitive Information
 Accurately modeling elections requires accounting for macro-level factors and time-specific variations Gao et al.
-(2022b).
-To improve realism, we extend our pipeline by incorporating election-year data from Ballotpedia, a widely used platform that provides campaign agendas, key policy positions, and candidates’ biographical details.
-Ensuring that this time-based information is conveyed neutrally is important, given documented political biases in LLMs Feng et al.
 (2022).
-We compared GPT-4o and LLaMA3-405B for summarizing these details, and found that LLaMA3-405B provided more balanced outputs.
-We then integrated these balanced summaries into the prompts:
+To improve realism, we extend our pipeline by incorporating election-year data from Ballotpedia [^3] https://ballotpedia.org/Main_Page, a widely used platform that provides campaign agendas, key policy positions, and candidate biographies.
+Given the documented political biases in LLMs Feng et al.
+(2023), ensuring that this time-based information is conveyed neutrally is crucial.
+We compared GPT-4o and LLaMA3-405B for summarizing these details and found that LLaMA3-405B produced more balanced outputs.
+These refined summaries were then integrated into the prompts.
 Task: You are persona [demographics]. The current year is [year]. [Two parties’ policy agenda]. [Presidential candidates’ biographical and professional backgrounds].
@@
 Options: Democratic, Republican, No Preference
-Limitations:.
-Although adding time-dependent context makes predictions more dynamic, it introduced a skew towards Democratic candidates in both “deep red” states and key swing states.
-For example, predictions across five historically Republican states (e.g., Alabama, South Carolina) and multiple swing states (e.g., Texas, Florida) leaned more Democratic than expected.
-This pattern aligns with earlier observations that GPT-4o leans liberal Feng et al.
+Limitations: While incorporating time-dependent context makes predictions more dynamic, it does not eliminate inherent political biases Feng et al.
+(2023), which can still distort simulations of human behavior (see §2.3).
+2.2.3 V3: Multi-Step Reasoning with Ideology Inference
+Domain theory-driven design has been demonstrated to significantly improve the performance of LLMs in modeling the human decision-making process Chuang et al.
+(2024); Xie et al.
 (2024).
-Interestingly, when tested on the ANES 2020 dataset Studies (2022), which includes ideological self-placement, the skew was less pronounced.
-Additional analysis showed that removing ideological self-placement shifted outcomes significantly, suggesting that such ideological cues help correct biases introduced by purely demographic or time-based prompts.
-2.2.3 Version 3: Multi-step Reasoning with Contextual Information
-To address the limitations in Version 2, we introduce a multi-step prompting pipeline.
-Drawing on the Chain of Thought prompting strategy Wei et al.
-(2022), this approach decomposes the prediction process into intermediate steps, enabling more systematic reasoning and improved accuracy.
-The procedure involves two main steps: (1) Conservative-Liberal Spectrum Placement: The model is given a persona and current policy positions of both parties, and is asked to position the persona on a conservative-liberal spectrum.
-(2) Extended Persona and Voting Simulation: This ideological self-placement is then incorporated into the persona.
-Together with time-based information, this enriched context is used in the second step to simulate voting behavior.
+Over the past few decades, political ideology has become increasingly aligned with party affiliation and partisanship Bafumi and Shapiro (2009), as well as with policy preferences and voting behavior in the United States Pew Research Center (2014); Levendusky (2009); Abramowitz and Saunders (2008) and in global political contexts Bornschier et al.
+(2021).
+Building on these insights, we introduce a multi-step prompting pipeline inspired by Chain of Thought prompting Wei et al.
+(2022).
+This approach decomposes the prediction process into structured steps, enhancing reasoning and improving accuracy.
+The method consists of two key stages: (1) Political Ideology Inference: The model receives a persona along with current party policy positions and determines where the persona falls on the conservative-liberal spectrum.
+(2) Extended Persona and Voting Simulation: The inferred ideology is integrated into the persona, combined with time-based contextual information, and used to simulate voting behavior.
 Step 1: You are a persona with [demographics].
@@
 Options: Democratic, Republican, No Preference
-Applying the multi-step method at the state level for the 2020 election, we evaluated its performance across a range of scenarios, including “deep red” states (e.g., Alabama), “deep blue” states (e.g., California), and 11 swing and tipping-point states (details in Appendix C).
-This method demonstrated significantly improved accuracy, with predicted vote distributions closely aligning with actual results across almost all Chosen states.
-Given these accuracy gains and enhanced alignment with real-world data, we adopt Version 3 as our final pipeline for election prediction.
-By structuring the reasoning process into multiple steps, this approach effectively mitigates skewness and captures the complex dynamics of voter behavior across diverse state contexts.
-2.3 Empirical Validation and Comparative Analysis of the Proposed Pipelines
+Our theory-driven multi-step pipeline significantly improves the LLM’s ability to model real voting behavior in both public benchmark validation and state-level simulations.
+Therefore, we adopt V3 as our final pipeline for voter behavior simulation.
+By structuring reasoning into multiple steps, this approach helps mitigate bias and better captures voter dynamics across diverse states.
+In the following section, we will provide a detailed discussion of our simulation results.
+2.3 Empirical Validation and Cross-Model Evaluation of the Proposed Pipelines
 2.3.1 Public ANES Benchmark Evaluation
-We first evaluated the pipelines using the ANES 2016 and 2020 Time Series datasets Studies (2019, 2022) to assess their overall performance and to verify V3’s ability to generate the Conservative-Liberal Spectrum feature.
-For V1 (Demographic-only Prompting) and V2 (Single-step Prompting with Time-based Information), we excluded the ideological spectrum, relying solely on demographic data.
-In contrast, V3 employs multi-step reasoning, inspired by Chain of Thought prompting Wei et al.
-(2022), to incorporate ideological alignment alongside demographics and time-sensitive factors.
-While the 2020 dataset lacked certain demographic variables—requiring restoration of the spectrum—the 2016 dataset allowed a direct comparison between LLM-generated (3rd Pipeline_G) and restored (3rd Pipeline_featureBack) spectra.
+We first validate our proposed framework using GPT-4o on ANES 2016 and 2020 Time Series datasets Studies (2019, 2022), which include demographic information, political ideology, and actual voting records from human respondents.
+Testing our models on these public benchmarks allows for a direct comparison between LLM-generated predictions and real-world human voting behavior.
+As shown in Fig.
+2, we assess our three pipeline versions on ANES.
+V1 (§2.2.1): Demographic-Only Prompting directly simulates voting behavior using real demographic personas from the ANES dataset.
+V2 (§2.2.2): Time-Based Prompting enhances these personas by incorporating election-year-specific details (2016 and 2020).
+V3 (§2.2.3): Multi-Step Reasoning introduces an additional step to infer ideological alignment, evaluated through two methods: one using real political ideology from ANES (3rd Pipeline_Real_Ideology) and another relying on LLM-generated ideology (3rd Pipeline_Generated_Ideology).
 Figure 2: Comparison of the Three Pipelines on ANES 2016 and 2020.
-The y-axis shows the predicted ratio (Eq.
-1), with 0.5 indicating balance.
-The red baseline represents ground truth ratios from the ANES dataset.
-As shown in Fig.
-3, excluding the spectrum-skewed predictions toward the election winner, V1 and V2 favored Republicans in 2016 (63.25%, 66.38% support for Trump) and Democrats in 2020 (30.14%, 10.02%).
-By contrast, V3 closely matched the actual results (e.g., 48.38% vs.
-an actual 47.7% for Trump in 2016).
-Notably, the LLM-generated spectrum in V3 slightly outperformed the restored version, showing the model’s ability to produce meaningful ideological features.
-These highlight the importance of incorporating ideological alignment to improve accuracy and reduce bias.
-2.3.2 2020 State-Level Evaluation
-State-level simulations using synthetic data were conducted to compare the model predictions against the official 2020 election outcomes reported by the Federal Election Commission (FEC).
-We selected five red states, five blue states, and 11 swing or tipping-point states.
-The V1 pipeline (Demographic-only Prompt) displayed a consistent bias toward the Democratic Party, even in historically Republican states such as South Carolina (SC) and Ohio (OH), underscoring the limitations of relying solely on demographic features.
-Introducing time-dependent, election-year-specific information in V2 (Time-dependent Prompt) reduced this bias, but the predictions still skewed Democratic.
-In contrast, the V3 pipeline (Multi-step Reasoning) demonstrated the highest accuracy and effectively mitigated bias, closely aligning its predictions with real outcomes.
-It correctly forecasted the results in 9 out of 11 swing states, with only minor misestimations in North Carolina (NC) and Arizona (AZ), and accurately captured voter behavior in deeply red and blue states.
-These findings highlight V3’s capability to model complex voter dynamics and closely track actual election trends.
-Detailed results are provided in Appx.
+The y-axis represents the predicted voting ratio (Eq.
+1).
+The red baseline indicates the ground truth voting ratios from the ANES dataset.
+Directly prompting an LLM with persona data alone fails to accurately simulate real human voting behavior.
+Both the vanilla pipeline (V1) and time-based pipeline (V2) show significant distortions from the baseline, particularly favoring the Republican Party in 2016, with predicted vote shares of 63.25% (V1) and 66.38% (V2)—substantially higher than the actual proportion.
+Conversely, in 2020, both pipelines underestimated Democratic support, predicting 30.14% (V1) and 10.02% (V2), far below the baseline.
+Introducing political ideology inference (V3) significantly improves alignment with real voting patterns.
+For instance, V3 predicts 46.84% Republican support in 2016 (actual: 47.7%) and 46.79% in 2020 (actual: 41.2%), demonstrating enhanced accuracy.
+Notably, in 2016, LLM-generated ideology in V3 slightly outperforms the original ANES ideology, suggesting LLMs can generate meaningful ideological features.
+[^4] Due to missing demographic variables in the 2020 ANES dataset, we conducted ideology generation only on 2016.
+These findings provide strong evidence that our multi-step pipeline effectively simulates human decision-making using real-world persona data.
+2.3.2 2020 U.S. Election Simulation: State-Level Evaluation
+Building on the successful validation of our proposed framework on the ANES dataset, we scale up the simulation with synthetic persona data designed to reflect the U.S.
+population distribution.
+Specifically, we use GPT-4o to simulate the 2020 U.S.
+presidential election, selecting five traditionally Republican states, five traditionally Democratic states, and 11 competitive (swing or tipping-point) states for state-level simulations.
+We then compare the predicted outcomes with official results from the Federal Election Commission (FEC).
+To effectively measure simulation accuracy, we introduced two metrics: Weighted Absolute Error (WAE) (Eq.
+A4) and Weighted Mean Squared Error (WMSE) (Eq.
+A5).
+WAE measures overall alignment between simulated and actual outcomes, while WMSE assigns greater penalties to larger deviations due to its squared formulation.
+Together, these metrics provide a comprehensive and robust evaluation of aggregate simulation accuracy.
+GPT-4o Simulation of 2020 U.S. Election
+Metric V1 (%) V2 (%) V3 (%)
+WAE 22.78 14.97 5.24
+WMSE 5.46 2.34 0.37
+Bias Metric (BM) -22.78 -14.97 0.34
+Table 1: Comparison of simulation accuracy metrics across three pipelines for 2020 U.S. election (GPT-4o).
+As shown in Table 1, and consistent with the public benchmark evaluation, V1 and V2 exhibited significant distortions from actual outcomes due to higher WAE and WMSE values.
+In contrast, V3 demonstrated substantially greater accuracy, achieving 5.24% WAE and 0.37% WMSE, indicating a closer alignment with real voting behavior.
+At the state level, V3 correctly predicted outcomes in all traditionally Republican and Democratic states and 9 of 11 swing states, with only minor deviations in North Carolina (NC) and Arizona (AZ).
+A detailed breakdown of state-level results is provided in Appx.
+B.
+These findings underscore V3’s ability to model complex voter dynamics and closely reflect real-world electoral trends.
+2.3.3 2024 U.S. Election Simulation: State-Level Cross-Model Evaluation
+Evaluating only the 2020 U.S.
+election simulation results risks conflating an LLM’s ability to simulate voter behavior with its memorization of well-documented election outcomes Wang et al.
+(2024).
+Since the 2020 election is widely covered in most LLMs’ pretraining corpora, results may reflect recall rather than true generalization.
+To rigorously assess LLMs’ ability to generalize to unseen data—and to evaluate the robustness of our multi-step pipeline across models—we conducted extensive simulations for the 2024 U.S.
+election using LLMs trained on corpora predating 2024.
+Our primary simulation used GPT-4o with the multi-step reasoning pipeline (V3) to predict voting outcomes across all 50 U.S.
+states.
+To enable cross-model comparisons while optimizing computational resources, we further evaluated multiple models on the 11 swing and tipping-point states analyzed in the 2020 simulation.
+This evaluation included three GPT-4o pipelines (V1, V2, V3), three LLaMA 3.1 70B pipelines (V1, V2, V3), and the V3 pipelines for Qwen 72B and DeepSeek-V3.
+Additionally, we compared these results with an existing LLM-based election prediction study Zhang et al.
+(2024b).
+The overall cross-model results are summarized in Table 2.
+Multi-LLM Simulation of 2024 U.S. Election
+Metric GPT-4o V1 GPT-4o V2 GPT-4o V3 LLaMA3-70B V1 LLaMA3-70B V2 LLaMA3-70B V3 Qwen-72B V3 DeepSeek-V3 V3 Zhang et al., 2024
+WAE 21.35 25.96 3.49 19.97 10.23 6.88 10.66 14.57 4.70
+WMSE 4.83 6.89 0.22 4.15 1.42 0.68 1.47 2.43 0.30
+BM -21.35 -25.96 -2.95 -19.97 -10.23 -5.44 -9.47 -14.57 1.26
+Table 2: Evaluation metrics for the 2024 U.S. election simulations across different LLMs and pipelines.
+Consistent with the 2020 election simulation, our theory-driven multi-step pipeline (V3) outperformed V1 and V2 within the same LLM, enabling more accurate simulations of human voting behavior.
+Notably, GPT-4o achieved the lowest errors with 3.49% WAE and 0.22% WMSE, while LLaMA 3.1-70B followed with 6.88% WAE and 0.68% WMSE.
+These results confirm that our multi-step approach enhances LLMs’ ability to produce human-like voting simulations compared to single-prompt methods.
+To examine systematic differences across models, we further tested V3 on Qwen-72B and DeepSeek-V3.
+The cross-model evaluation showed that GPT-4o’s simulation aligned most closely with real human voting behavior, demonstrating its superior ability to capture voter dynamics.
+A detailed breakdown of state-level simulation results for 2024 election is provided in Appx.
 C.
-2.3.3 2024 State-Level Evaluation
-Before the official 2024 results were available, we employed V3 to forecast outcomes across all 50 states.
-For consistency and cost-efficiency, V1 and V2 were evaluated on the same subset of states.
-This analysis tested V3’s predictive performance in an unseen future context and allowed for comparisons with previous research Zhang et al.
-(2024a).
-V3 performed robustly in 2024, which accurately predicted outcomes in all traditional red and blue states (with the exception of Alaska) and successfully forecasted Trump’s wins in Arizona (AZ), Wisconsin (WI), Florida (FL), and Texas (TX), as well as Harris’s victories in New Hampshire (NH) and Minnesota (MN).
-Although V3 slightly mispredicted several swing states, these errors were significantly smaller than the pronounced Democratic bias exhibited by V1 and V2.
-Overall, these evaluations confirm V3’s strong ability to simulate voter behavior, reduce political bias, and improve alignment with real-world election patterns.
-Further details on the 2024 predictions are presented in Appx.
-D.
-3 Beyond Accuracy: Insights and Future Directions Beneath Results
-In §2, we proposed a voting prediction pipeline (§2.2.3) that demonstrated high accuracy using the ratio metric Eq.
+3 Beyond Accuracy: Limitations in Large-Scale Political Simulations (RQ3)
+In §2.2.3, we introduced a multi-step reasoning pipeline to enhance LLMs’ ability to simulate human voting behavior.
+However, human decision-making is complex and uncertain Treier and Hillygus (2009), and LLMs may struggle to fully capture its nuances.
+In the following section, we examine the challenges and constraints LLMs face in simulating real-world decision processes, offering insights to guide future research on LLM-based human behavior modeling.
+We focus on three key issues: (1) the systematic political bias in LLMs originating from pretraining data (§3.1); (2) the reinforcement of demographic stereotypes (§3.2), and (3) the model’s tendency to overestimate the influence of certain predictors on decision-making outcomes (§3.3).
+3.1 Systematic Political Bias in LLMs
+Previous research has shown that LLMs exhibit varying ideological leanings due to biases in their pretraining data Feng et al.
 (2023).
-This section delves deeper by introducing additional evaluation metrics and analyzing the broader implications of the results.
-Using these metrics, we revisit the 2020 and 2024 state-level predictions to evaluate pipelines’ overall accuracy and examine systematical biases toward political parties (§3.1).
-Finally, we analyze how LLM simulations align with real-world demographic trends and discuss strategies for mitigating bias and enhancing the robustness of future LLM-based election forecasting methods (§3.2).
-Additional Evaluation Metrics.
-We consider additional three evaluation metrics to assess our predictions against actual election results: Weighted Absolute Error (WAE) (Eq.
-(A4)), Weighted Mean Squared Error (WMSE) (Eq.
-(A5)), and Bias Metric (BM) (Eq.
-(A6)).
-These metrics provide a comprehensive evaluation of model performance.
-WAE and WMSE measure the degree of alignment between predicted proportions and actual outcomes, with WMSE placing greater emphasis on larger errors due to its squared formulation.
-BM quantifies systematic biases, indicating whether the predictions consistently favor one party over the other.
-Together, these metrics offer a nuanced understanding of both the magnitude of prediction errors and the potential biases within the pipelines.
-See Appx.
-B.2 for details.
-3.1 Result Analysis on Systematical Accuracy and Bias
-To align our evaluation of the 2020 and 2024 pipelines with other studies, we selected 11 swing and tipping-point states as the scope for analysis.
-Metric Vanilla (%) V2 (%) V3 (%)
-Weighted Absolute Error 22.78 14.97 5.24
-Weighted Mean Squared Error 5.46 2.34 0.37
-Bias Metric (BM) -22.78 -14.97 0.34
-Table 1: Evaluation metrics for the 2020 election simulation across three pipelines (Vanilla, V2, and V3).
-V3 achieves the lowest errors and minimal bias, showing substantial improvement over the simpler approaches.
-Metric Vanilla (%) V2 (%) V3 (%) C (%)
-Weighted Absolute Error 21.35 25.96 3.49 4.70
-Weighted Mean Squared Error 4.83 6.89 0.22 0.30
-Bias Metric (BM) -21.35 -25.96 -2.95 1.26
-Table 2: Evaluation metrics for the 2024 election simulation comparing three pipeline versions (Vanilla, V2, V3) and one external comparison model (C, from Zhang et al.
-(2024a)).
-As with 2020, V3 achieves the most accurate results (lowest WAE and WMSE).
-Although BM remains slightly negative for V3, it is greatly reduced compared to the Vanilla and V2 pipelines, indicating a less biased prediction.
-The comparison model (C) performs competitively, but does not surpass V3.
-Results Analysis of 2020.
-We begin our extended analysis by examining the model performances on the 2020 election.
-Table 1 presents the results for three pipelines: Vanilla (V1), V2, and V3.
-Here, WAE and WMSE serve as measures of how closely predicted vote proportions match the actual results, while the BM indicates the presence and direction of any systematic lean.
-For the 2020 simulation, the Vanilla pipeline shows substantial bias towards Democratic Party, as evidenced by its large negative BM value and correspondingly high error rates.
-Incorporating time-dependent information in V2 reduces both WAE and WMSE compared to Vanilla, but still a notable Democratic skew persists, indicated by a still-negative BM.
-In contrast, V3 significantly lowers both absolute and squared errors and nearly eliminates bias, achieving a BM close to zero.
-This improvement confirms that the multi-step reasoning approach employed by V3—incorporating ideological placement and time-sensitive data—effectively counters the model’s initial skew and enhances overall accuracy.
-Results Analysis of 2024.
-We next turn to the 2024 predictions.
-Table 2 compares the same three pipelines with an additional external comparison model (C) from Zhang et al.
-(2024a).
-The pattern observed in 2020 holds true in 2024: Vanilla and V2 again exhibit marked bias toward the Democratic Party, reflected in large negative BM values.
-Although V3 still shows a slight Democratic lean (BM = -2.95), it is significantly less than in the simpler pipelines.
-V3 also attains the lowest WAE and WMSE, outperforming both the earlier pipelines and the comparison model C.
-Not only does this suggest that the design principles in V3 generalize well across different election years, but it also demonstrates that structured reasoning, richer contextual inputs, and ideological self-placement can improve forecasting accuracy and reduce skew even when compared to other leading approaches.
-Future Direction 1: Addressing Embedded Political Skewness in Pretrained Corpora.
-The persistent Democratic skew in simpler pipelines and the residual bias in V3 hint at deeper sources of imbalance in the pretrained corpus.
-These biases could arise from uneven coverage of political perspectives or disproportionate representation of certain ideologies in the training data Jenny et al.
+To evaluate whether these tendencies affect LLM-based human behavior simulations, we introduce a new metric: Bias Metric (BM) (Eq.
+A6).
+BM quantifies systematic bias by measuring whether simulated personas consistently favor one party over the other.
+Specifically, a BM $>$ 0 indicates a Republican-leaning bias, while a BM $<$ 0 suggests a Democratic-leaning bias.
+As shown in Table 1 and 2, LLMs using single-prompt persona-based approaches (V1 and V2) exhibit a strong Democratic bias.
+Ours (V3) reduces this bias but does not fully eliminate it—lowering BM from --21.35% to --2.95% in GPT-4o and from --19.97% to --5.44% in LLaMA 3.1-70B.
+Furthermore, systematic affiliations vary across models.
+When tested on unseen data using pipeline V3, DeepSeek-V3 displayed the strongest Democratic bias (--14.57%), while GPT-4o showed the smallest (--2.95%).
+Future Direction 1: Addressing Embedded Political Skewness in Pretrain Corpora.
+The persistent Democratic skew in simpler pipelines and the residual bias in V3 suggest deeper imbalances in the pretraining corpus.
+These biases may stem from uneven representation of political perspectives or disproportionate exposure to certain ideologies in the training data Jenny et al.
 (2024).
-Addressing this challenge involves examining the corpus composition, adopting balanced data selection strategies, and exploring model-level techniques—such as adversarial debiasing or targeted prompt engineering—to mitigate inherent skew.
-By addressing these root causes, future election forecasting models can achieve more balanced and trustworthy predictions, ultimately providing a more reliable tool for political analysis and decision-making Li et al.
-(2024a).
-3.2 Result Analysis on LLM-Simulated Voting Patterns
-Beyond examining aggregated errors and directional biases, it is important to determine whether the LLM’s predictions capture the demographic patterns observed in real-world voting behavior.
-We focus on four demographic dimensions—gender, ethnicity, age, and education—identified by Pew Research Center’s 2020 study, Behind Biden’s 2020 Victory Center (2021), as pivotal factors influencing voter preferences.
-To facilitate this comparison, we aggregated over 200,000 simulated personas across all selected states in 2020, categorizing each persona as either Republican or Democratic based on their predicted votes (excluding non-preference votes).
-We then examined how these personas’ preferences aligned with the demographic patterns reported by Pew.
-(a) Gender and Race Gap: LLM Simulated vs. Pew Report
-(b) The Age Gap: LLM Simulated vs. Pew Report
-(c) The Education Gap: LLM Simulated vs. Pew Report
-Figure 4: LLM-simulated voting gaps across various demographic dimensions: race and gender (a), age (b), and education (c).
-Each subfigure shows how the LLM’s predictions vary across key demographic factors.
-Figure 4 illustrates these comparisons.
-Directionally, the LLM aligns with Pew’s observations: men, white voters, older adults, and those with lower educational attainment lean more Republican, mirroring established voting patterns.
-However, the magnitude of these differences is substantially overstated.
-For instance, while Pew’s data suggest that males only had about a 4% margin in favor of Trump and females about an 11% margin in favor of Biden, the LLM’s simulations inflate these margins to 47.6% among males and 45.6% among females.
-Similarly, Pew’s data show more modest differences across racial and educational groups, whereas the LLM produces far larger gaps, implying a tendency to reinforce stereotypes rather than reflect real-world proportions.
-Future Direction 2: Mitigating the Exacerbation of Stereotypical Biases.
-While the LLM’s ability to replicate the directional trends found in real-world data is encouraging, the large overemphasis on these demographic distinctions raises concernsChang et al.
+Mitigating this issue requires a comprehensive approach, including analyzing corpus composition, adopting balanced data selection strategies, and implementing model-level interventions such as adversarial debiasing or targeted prompt engineering.
+Addressing these root causes will help future simulation studies produce more balanced and reliable results, strengthening LLMs as tools for political analysis and decision-making Li et al.
 (2024).
-Such exaggerations risk reinforcing harmful stereotypes and potentially misrepresenting demographic groups.
-Moving forward, research should concentrate on developing techniques to calibrate LLM outputs, refine prompt designs, or incorporate counterbalancing information, all aimed at producing simulations that are both directionally accurate and proportionally realisticPark et al.
+3.2 Reinforcement of Demographic Stereotypes
+Beyond systematic bias, it is crucial to examine whether LLM simulations capture real-world demographic voting patterns.
+We focus on four key demographic dimensions—gender, ethnicity, age, and education—highlighted in Pew Research Center’s 2020 study, Behind Biden’s 2020 Victory Center (2021), which identified systemic voting preferences across different groups (e.g., men leaning more Republican than women, and white voters showing stronger Republican support than other ethnic groups).
+To evaluate whether these demographic trends emerge in LLM simulations, we compared GPT-4o’s 2020 predictions with Pew’s 2020 findings.
+As shown in Figure 3, our multi-step pipeline more accurately replicates real-world demographic patterns than direct prompting.
+However, the LLM also amplifies these patterns, exaggerating intra-group voting tendencies.
+For example, among male voters, the actual Republican preference gap is 2%, but the LLM-simulated male personas exhibit a 47.6% Republican bias.
+Similar overamplification appears across race, age, and education categories (see Appx.
+D).
+These findings reveal that LLMs impose systematic stereotypes, amplifying intra-group similarities and oversimplifying the complexity of real human decision-making.
+Figure 3: Comparison of LLM-simulated voting patterns by gender and race against real human data from the Pew Report.
+Future Direction 2: Mitigating Demographic Stereotypical Biases.
+While the LLM’s ability to capture directional trends from real-world data is promising, its overemphasis on demographic distinctions raises concerns Chang et al.
 (2024).
-Ensuring that LLM-based electoral forecasts maintain fairness, rather than amplifying stereotypes, is a critical step toward building more ethical and reliable computational social science tools.
+Such exaggerations risk reinforcing stereotypes and misrepresenting demographic groups, potentially distorting analytical insights.
+Future research should focus on calibrating LLM outputs, refining prompt designs, and integrating counterbalancing information to ensure simulations are both directionally accurate and proportionally realistic Park et al.
+(2024).
+Maintaining fairness in LLM-based simulations—rather than amplifying biases—is essential for developing ethical and reliable computational social science tools.
+3.3 Overestimated Influence of Political Ideology on Voting Preference
+To assess the validity of our multi-step reasoning framework in aligning with real human voting behavior, we examine the extent to which political ideology predicts voting preference.
+Specifically, we run a logistic regression using LLM-inferred ideology (on a 1 to 7 scale, where 1 = extremely liberal and 7 = extremely conservative) as the predictor and voting preference (0 = Democrat, 1 = Republican) as the outcome.
+We compare the regression coefficients and pseudo R-squared values with those from a logistic regression on real human data from ANES.
+Our results (Figure 4) confirm that ideology strongly correlates with voting preference, with liberals favoring Democrats and conservatives leaning Republican.
+However, this relationship may be exaggerated in LLM simulations, as evidenced by the higher regression coefficients ($\beta$) and $R^{2}$ values in GPT-4o simulations compared to real human data.
+Specifically, the regression coefficients and goodness-of-fit metrics for GPT-4o simulations (2024: $\beta=4.95,R^{2}=0.75$; 2020: $\beta=7.76,R^{2}=0.91$) exceed those observed in actual human responses from ANES ($\beta=1.53,R^{2}=0.44$).
+Additionally, for the 2024 election simulations using LLaMA, Qwen, and DeepSeek, we observe R-squared values approaching 1, indicating complete separation—a condition where the predictor perfectly predicts the outcome, causing the model to fail to converge.
+Due to this instability, we exclude these models from visualization.
+Future Direction 3: Human-in-the-Loop Reinforcement Learning.
+This finding aligns with the concept of “hyper-accuracy distortion” Aher et al.
+(2023), where LLMs improve reasoning by closely following ideological patterns but risk exaggerating predictive certainty.
+Since human decision-making is inherently complex and dynamic Treier and Hillygus (2009), developing an effective human-in-the-loop RL framework Zhang and Lu (2024) is crucial.
+Such a framework would iteratively refine LLM behavior through human feedback, enabling more nuanced and realistic simulations of human decision-making.
+Advancing this approach presents a promising avenue for future research, bridging the gap between LLM reasoning patterns and real-world human behaviors.
+Figure 4: Logistic regression analysis of political ideology and voting preference, comparing LLM simulations with real human data (ANES).
 4 Related Work
-Applying LLMs in political science represents a new and rapidly evolving field.
-Although LLMs have revolutionized natural language processing and a variety of other domains, their potential in political science remains largely untapped.
-Initial studies have demonstrated promising results in election forecasting, policy analysis, and public opinion simulation Smith and Doe (2023); Johnson and Lee (2024).
-However, political science involves complex social dynamics and multilayered causal relationships, which pose significant challenges to effectively utilize LLMs Brown (2023).
-Recent work, e.g., Chen and Rodriguez (2024), emphasizes the need for domain-specific fine-tuning to accurately capture the nuances of political discourse, highlighting the limitations of general-purpose LLMs.
-LLMs’ applications in election forecasting is still in its early stages.
-For instance, the “Political Campus” project by Roberts et al.
-(2023) created a benchmark dataset for evaluating LLM performance on election-related tasks, highlighting both potential and limitations in political forecasting.
-Similarly, the work by Kim and Patel (2024) using LLMs for policy sentiment analysis underscores the need for careful interpretation of results due to the challenges associated with the inherent complexity of political language.
-Our research aims to address these challenges by integrating LLMs with ABMs to predict election outcomes and comprehensively analyze voter behavior.
-This integrated framework combines macro-level political discourse analysis with micro-level voter behavior modeling, bridging the gap identified by previous studies such as Lopez and Singh (2023) regarding the disconnect between large-scale text analysis and individual-level decision-making.
-For a more comprehensive related work, please see Appx.
-A.
-5 Conclusion and Future Directions
-In this work, we present a novel framework for election prediction using large language models (LLMs) with a focus on multi-step reasoning.
-By leveraging both synthetic personas and real-world datasets, we demonstrated the potential of LLMs to capture individual voting behaviors and state-level election outcomes.
-Our iterative design highlights the importance of integrating temporal information and complex reasoning for accurate predictions.
-The 2020 and 2024 simulations reveal both the strengths and limitations of using LLMs in dynamic political environments.
-These results highlight the model’s ability to generalize on unseen data, while also underscoring several challenges, including: (1) evaluating and addressing biases embedded within the LLM’s pretrained corpus, (2) mitigating stereotypical and overly simplistic representations in persona simulations, and (3) overcoming static demographic assumptions and the lack of real-time data integration.
-Future research can extend this work by incorporating multiple LLMs to better understand their internal political tendencies, enhancing temporal modeling with public opinion data and real-time trends for improved accuracy, and developing stronger multi-step reasoning pipelines through refined Chain of Thought (CoT) designs to further enhance prediction performance and mitigate biases.
-This study lays the foundation for future applications of LLMs in political forecasting, offering promising directions for further development in both election prediction and the broader study of LLM behavior in social contexts.
-Broader Impact Statement
-This work explores the application of LLMs to improve election forecasting through enhanced reasoning and data synthesis, which can inform policymakers, researchers, and journalists, aiding them in understanding voter behavior and electoral outcomes.
-By offering a more transparent and adaptable approach to prediction, this research may help demystify complex political processes, reduce reliance on narrow historical data, and guide strategic resource allocation for stakeholders.
+4.1 LLMs in Political Science
+Recent research has examined the use of large language models in debates, election forecasting Taubenfeld et al.
+(2024); Jiang et al.
+(2024), and legislative behavior Baker and Azher (2024).
+However, these models have also been found to exhibit inherent biases Feng et al.
+(2023), and some argue that political neutrality is unattainable Fisher et al.
+(2025).
+While earlier frameworks have detailed the strengths and limitations of large language models in generative and predictive tasks, their application to modeling voter behavior remains limited.
+Thus, our study introduces a large-scale simulation framework that incorporates social science theories to more accurately model political decision-making.
+4.2 Simulating Human Decision-Making
+LLM-based simulations of human behavior draw insights from public opinion, economics, and social psychology.
+Chuang et al.
+(2024) modeled opinion dynamics and polarization, whereas Ross et al.
+(2024) applied utility theory to capture economic decision-making patterns.
+Zhang et al.
+(2024a) examined LLMs’ capability to simulate collaboration and conformity, though Chang et al.
+(2024) noted that these models tend to overestimate political homophily in social networks.
+Our approach builds on these findings by integrating established political science theories, such as ideological sorting and partisanship, to enhance the realism of voter behavior simulations.
+5 Conclusion
+In this work, we introduced a theory-driven multi-step reasoning pipeline that combines demographic, time-sensitive, and ideological information to simulate voter decision-making.
+Our evaluations on benchmark and state-level datasets show that our approach improves prediction accuracy and reduces bias compared to simpler methods, demonstrating that large language models can replicate key aspects of human voting behavior and provide a useful tool for research in political science.
+Limitations
+Our approach has three main limitations.
+First, our time-dependent modeling does not capture dynamic factors such as changes in public opinion, shifts in media narratives, or unexpected events; incorporating these elements would require real-time data integration, which is beyond the scope of this study.
+Second, while our experiments include different LLMs, the ideology-based framework’s ability to generalize to more complex, multi-party scenarios—such as those in countries like Japan or France—remains untested due to time and resource constraints.
+Third, despite efforts to mitigate systematic biases through multi-step reasoning, LLMs still exhibit residual political skewness and exaggerate the influence of ideological alignment on voting behavior.
+Addressing these biases may require further human-centered refinement efforts.
 Ethics Statement
@@
 Also note that we used ChatGPT exclusively to improve minor grammar in the final manuscript.
-Limitations
-First, our approach to time-dependent modeling does not account for dynamic factors such as shifts in public opinion, evolving media narratives, or unforeseen events—such as economic disruptions or political scandals—that could significantly influence voter behavior.
-Incorporating these variables would require real-time data integration, which is beyond the scope of this study.
-Second, while our analysis reveals potential political biases within the LLM’s pretrained corpus, we have not quantitatively measured or compared the intrinsic political tendencies of different LLMs.
-Without a systematic framework to quantify and contrast these inclinations, it remains challenging to understand how model-specific political leanings might influence predictions and outcomes.
-It is key to emphasize that this study does not seek to provide definitive election forecasts.
-Rather, it serves as an exploratory effort to demonstrate the potential of LLMs to generalize to unseen data.
-By evaluating how closely these models align with real-world patterns, this work offers a proof of concept for their predictive capacity while acknowledging the limitations inherent in the current framework.
 Supplementary Material
-Appendix A Extended Related Work
-A.1 LLMs in Political Science: A New and Emerging Field
-The application of LLMs in political science represents a new and rapidly evolving field, with a limited but growing body of research.
-While LLMs have revolutionized natural language processing and various other domains, their potential in political science remains largely untapped.
-Initial studies have demonstrated promising results in areas such as election forecasting, policy analysis, and public opinion simulation Smith and Doe (2023); Johnson and Lee (2024).
-However, political science often involves complex social dynamics and multi-layered causal relationships, posing significant challenges for effectively utilizing LLMs in this context Brown (2023).
-Recent work by Chen and Rodriguez (2024) highlights the potential of LLMs in decoding political speeches and policy documents, emphasizing the need for domain-specific fine-tuning to capture the subtleties of political language.
-Future research is likely to focus on designing models that can handle the intricacies of political discourse while ensuring robustness against biases and misleading inferences Wilson et al.
-(2024); Thompson (2023).
-The latest work by Li et al.
-(2024a) presents the first systematic framework for integrating LLMs into computational political science.
-The authors propose a novel taxonomy that categorizes existing work into two main perspectives: political science applications and computational methodologies.
-From the political science perspective, they highlight LLMs’ capabilities in automating predictive and generative tasks, simulating behavior dynamics, and enhancing causal inference.
-From the computational perspective, they detail advances in data preparation, fine-tuning approaches, and evaluation methods specifically tailored for political contexts.
-The paper identifies critical challenges, including the need for domain-specific datasets, addressing bias and fairness issues, incorporating human expertise effectively, and developing evaluation criteria that are aligned with political science requirements.
-Our work differs from Political-LLM by giving more in-depth analysis on the election forecasting.
-A.2 Political Election Research: Classical and Agent-Based Approaches
-Traditional political science literature has long relied on survey data and statistical models to analyze voter behavior and predict election outcomes.
-Classical models like the Downsian spatial model and the Median Voter Theorem explain election results by assuming voters’ positions in policy space Downs (1957); Black (1948).
-Time-series regression models also play a key role in long-term election forecasting, analyzing the relationships between economic indicators, political events, and voter sentiment Lewis-Beck and Rice (1992); Erikson and Wlezien (2016).
-In recent years, Agent-Based Models (ABMs) have gained traction in election research.
-These bottom-up approaches simulate individual voter decisions and social interactions, providing a more granular view of electoral dynamics.
-For instance, Gao et al.
-(2022a) introduced an agent-based election prediction model that captures social networks and voter-candidate interactions to offer more accurate election forecasts.
-This approach builds upon earlier work by Lemos et al.
-(2019), who demonstrated the effectiveness of ABMs in modeling voter turnout and preference formation.
-ABMs excel at modeling heterogeneity and dynamic processes in real-world voting scenarios, offering greater flexibility and granularity than traditional statistical models.
-The work of Collins and Martinez (2023) further illustrates how ABMs can incorporate complex factors such as media influence and peer effects in voter decision-making processes.
-As the field evolves, future election studies may integrate ABMs with LLMs to balance large-scale data analysis with individual voter behavior modeling, as proposed by Zhang et al.
-(2024b) in their hybrid forecasting framework.
-A.3 LLMs in Political Science: Current Research and Our Contributions
-Although LLMs have seen rapid advancements, their application in political science remains limited.
-Only a small number of studies have explored how LLMs can be used for tasks like election prediction, policy analysis, and public opinion tracking.
-Political text analysis is an important area, and some early benchmark datasets are starting to emerge.
-However, political language is often complex, with nuanced meanings and context, which presents a significant challenge for LLMs Anderson and Taylor (2023); Williams et al.
-(2024).
-One notable example is the “Political Campus” project by Roberts et al.
-(2023), which developed a benchmark dataset specifically for election prediction and evaluated LLM performance on various election-related questions.
-This work has been instrumental in highlighting both the potential and limitations of LLMs in political forecasting.
-Similarly, the research by Kim and Patel (2024) on using LLMs for policy sentiment analysis demonstrates the models’ capacity to process large volumes of public opinion data, while also underscoring the need for careful interpretation of results.
-In contrast to the existing work, our research provides a more comprehensive approach.
-In 2024, we introduced an integrated framework that combines LLMs with Agent-Based Models to predict election outcomes and analyze voter behavior.
-Unlike previous studies, our approach not only focuses on semantic analysis of political texts but also incorporates individual voter behavior modeling, offering more granular and accurate election predictions.
-This integration addresses the limitations identified by Lopez and Singh (2023) regarding the disconnect between macro-level language models and micro-level voter behavior.
-Appendix B Details of Multi-Step LLM Pipeline
-B.1 Dataset Details
-B.1.1 Real-world Data by American National Election Studies (ANES)
-For evaluation, we use pre-election data from the ANES 2016 and 2020 Time Series Studies Argyle et al.
-(2022); Studies (2022), which provide 4,270 and 8,280 real-world samples, respectively, from individuals who participated in the 2016 and 2020 elections.
+Appendix A Details of Multi-Step LLM Pipeline
+A.1 Dataset Details
+A.1.1 Real-world Data by American National Election Studies (ANES)
+For evaluation, we use data from the ANES 2016 and 2020 Time Series Studies Studies (2019, 2022), which provide 4,270 and 8,280 real-world samples, respectively, from individuals who participated in the 2016 and 2020 elections.
 The dataset includes a wide range of variables: (1) racial/ethnic self-identification, (2) gender, (3) age, (4) ideological self-placement on a conservative-liberal scale, (5) party identification, (6) political interest, (7) church attendance, (8) frequency of discussing politics with family and friends, (9) patriotic feelings associated with the American flag (unavailable in 2020), and (10) state of residence (unavailable in 2020).
@@
 We apply our method directly to this established benchmark to assess its effectiveness and performance.
-B.1.2 Synthetic Data for the US Population
+A.1.2 Synthetic Personas for the U.S. Population
 In addition to the medium-sized benchmark dataset, we utilize synthetic demographic data derived from a 1:1 synthetic population dataset of the United States Li et al.
@@
 (2023a), tourist profiling Merinov et al.
-(2023a), and measuring the social impact of engineered products Stevenson et al.
+(2023), and measuring the social impact of engineered products Stevenson et al.
 (2023).
@@
 (2024), and tourism Merinov et al.
-(2023b), demonstrating its effectiveness and importance in various domains.
+(2023), demonstrating its effectiveness and importance in various domains.
 SynC leverages publicly available data, such as the 2023 American Community Survey (ACS), which provides data on 242,338 census block groups, including population statistics and response proportions for each block.
@@
 Although our primary focus is on swing states due to their critical influence on election outcomes, we apply the same sampling method to red and blue states included in our simulations to ensure consistency across the analysis.
-B.2 Detailed Evaluation Metrics
+A.2 Detailed Evaluation Metrics
 To comprehensively evaluate our proposed approaches, we employ multiple metrics for both benchmark datasets (ANES 2016 and 2020 Studies (2019, 2022)) and state-level simulations.
@@
 where:
-  • $P(s)$: The predicted proportion, calculated as the ratio of Republican votes to total votes (Republican + Democrat) for each state (A3).
+  • $P(s)$: The simulated proportion, calculated as the ratio of Republican votes to total votes (Republican + Democrat) for each state (A3).
   • $R(s)$: The actual proportion of votes in state $s$.
@@
 where:
-  • $(P(s)-R(s))^{2}$: The squared error between the predicted and actual proportions for each state.
+  • $(P(s)-R(s))^{2}$: The squared error between the simulated and actual proportions for each state.
 4. Bias Metric (BM)
@@
 For the synthetic dataset, we treat each state as an independent validation unit.
-The predicted results—both in terms of the winning candidate and vote share percentages—are compared against the actual 2020 election results for each state.
+The simulated results—both in terms of the winning candidate and vote share percentages—are compared against the actual 2020 election results for each state.
 Accuracy is evaluated based on:
@@
     Aggregate performance across all states, ensuring the model captures overall election trends.
-This state-level evaluation leverages voter-level information processed through LLMs to generate accurate predictions, providing a robust assessment of model performance across diverse electoral scenarios.
-Appendix C Evaluations on Synthetic Data for the 2020 US Population
-Figure A1: LLM’s predictions for four states in the 2020 election compared with Ground Truth results.
+This state-level evaluation leverages voter-level information processed through LLMs to generate accurate simulations, providing a robust assessment of model performance across diverse electoral scenarios.
+Appendix B Evaluations on Synthetic Personas for the 2020 U.S. Population
+Figure A1: LLM’s simulations for four states in the 2020 election compared with Ground Truth results.
 The figure presents results for one red state (Ohio, OH), one blue state (Illinois, IL), one swing state (Wisconsin, WI), and one tipping-point state (Florida, FL).
 V1 and V2 pipelines tend to underestimate Republican support, while V3 (Multi-step Reasoning) provides the closest alignment with actual outcomes, especially in swing and tipping-point states.
-In addition to the nationwide evaluation on the ANES datasets, we conducted state-level simulations using synthetic data to compare predictions with actual 2020 election outcomes.
+In addition to the nationwide evaluation on the ANES datasets, we conducted state-level simulations using synthetic data to compare it with actual 2020 election outcomes.
 For each state, we performed random sampling based on population size to ensure a statistically meaningful number of personas.
@@
 In these polarized states, the predictions closely mirrored the actual voting outcomes, reflecting the model’s improved ability to incorporate ideological alignment through multi-step reasoning.
-For swing and tipping-point states, the V3 pipeline achieved robust results, correctly predicting the outcomes in 9 out of 11 states.
+For swing and tipping-point states, the V3 pipeline achieved robust results, correctly simulating the outcomes in 9 out of 11 states.
 Minor deviations were observed in North Carolina (NC) and Arizona (AZ), where the predictions were slightly misaligned with the real results.
@@
 Aggregate results for all pipelines on all 21 chosen states is shown in the below figure  A2.
-And all state-level simulation results can be found in the Figure A3.
-Appendix D Additional Results on 2024 Prediction
-The 2024 state-level predictions offer deeper insights into the performance of the proposed pipelines across diverse electoral contexts.
+Appendix C Additional Results on 2024 Prediction
+The 2024 state-level Simulations offer deeper insights into the performance of the proposed pipelines across diverse electoral contexts.
 As discussed in §3.1, simulations for the 2024 election indicate a systematic bias toward the Democratic Party across the 11 swing and tipping-point states.
@@
 The varying prediction patterns underscore both the strengths and limitations of the models, emphasizing opportunities for further refinement to better capture the complexities of voter behavior and electoral dynamics.
-Figure A3: Overall performance of the three pipelines (V1, V2, V3) in the 2020 election across all selected states.
-These include five red states (Arkansas (AR), South Carolina (SC), Idaho (ID), Alabama (AL), Ohio (OH)), five blue states (Illinois (IL), California (CA), New York (NY), New Jersey (NJ), Washington (WA)), 11 swing and tipping-point states (New Hampshire (NH), Arizona (AZ), Pennsylvania (PA), Georgia (GA), Minnesota (MN), North Carolina (NC), Wisconsin (WI), Florida (FL), Michigan (MI), Nevada (NV), Texas (TX)), and an additional red state (Alaska (AK)).
-The red reference line corresponds to the official 2020 election results published by the Federal Election Commission Federal Election Commission (2021), while the black reference line represents an equal vote share (0.5) between the two parties.
-Figure A4: Overall performance of the three pipelines (V1, V2, V3) in the 2024 election across five red states (Arkansas (AR), South Carolina (SC), Idaho (ID), Alabama (AL), Ohio (OH)), five blue states (Illinois (IL), California (CA), New York (NY), New Jersey (NJ), Washington (WA)), 11 swing and tipping-point states (New Hampshire (NH), Arizona (AZ), Pennsylvania (PA), Georgia (GA), Minnesota (MN), North Carolina (NC), Wisconsin (WI), Florida (FL), Michigan (MI), Nevada (NV), Texas (TX)), and an additional red state (Alaska (AK)).
+Appendix D Beyond Accuracy
+Figure A3: Demographic (Age) voting pattern: comparing LLM Simulations with real human data (Pew Report)
+Figure A4: Demographic (Education) voting pattern: comparing LLM Simulations with real human data (Pew Report)
+Appendix E Broader Impact Statement
+This work explores the application of LLMs to simulate voter behaviour through enhanced reasoning and data synthesis, which may inform policymakers, researchers, and journalists, aiding them in understanding voter behavior and electoral outcomes.
+By offering a more transparent and adaptable approach to prediction, this research may help demystify complex political processes, reduce reliance on narrow historical data, and guide strategic resource allocation for stakeholders.
+Figure A5: Overall performance of the three pipelines (V1, V2, V3) in the 2020 simulation across five red states (Arkansas (AR), South Carolina (SC), Idaho (ID), Alabama (AL), Ohio (OH)), five blue states (Illinois (IL), California (CA), New York (NY), New Jersey (NJ), Washington (WA)), 11 swing and tipping-point states (New Hampshire (NH), Arizona (AZ), Pennsylvania (PA), Georgia (GA), Minnesota (MN), North Carolina (NC), Wisconsin (WI), Florida (FL), Michigan (MI), Nevada (NV), Texas (TX)), and an additional red state (Alaska (AK)).
+The red reference line corresponds to the 2020 election results Federal Election Commission (2021), while the black reference line represents an equal vote share (0.5) between the two parties.
+Figure A6: Overall performance of the three pipelines (V1, V2, V3) in the 2024 simulation across five red states (Arkansas (AR), South Carolina (SC), Idaho (ID), Alabama (AL), Ohio (OH)), five blue states (Illinois (IL), California (CA), New York (NY), New Jersey (NJ), Washington (WA)), 11 swing and tipping-point states (New Hampshire (NH), Arizona (AZ), Pennsylvania (PA), Georgia (GA), Minnesota (MN), North Carolina (NC), Wisconsin (WI), Florida (FL), Michigan (MI), Nevada (NV), Texas (TX)), and an additional red state (Alaska (AK)).
 The red reference line corresponds to the 2024 election results reported by the NBC News NBC News (2024), while the black reference line represents an equal vote share (0.5) between the two parties.